### v1

# A solution to Banach conjecture

Abstract.
In this paper, we first prove that the origin-symmetric star body with ellipsoidal sections is an ellipsoid. Then we give a complete proof of Banach’s isometric subspace problem in finite dimensions through the John ellipsoid.
2010 Mathematics Subject Classification:
52A20, 52A22
1. Introduction
In this paper, we establish a positive answer to the following problem for finite $n$.
Problem 1.1.
Let $(V,\|\cdot\|)$ be a normed vector space (over $\mathbb{R}$) such that for some fixed $n$, $2\leq n\leq\mbox{dim}(V)$, all $n$-dimensional linear subspaces of $V$ are isometric. Is $\|\cdot\|$ neccessarily a Eulidean norm (i.e. an inner product one)?
This problem goes back to Banach in 1932 [2, Remarks on Chapter XII] and is often referred to as the ”Banach conjecture.” The conjecture asserts that the answer is always affirmative. It has been confirmed in many, though not all, dimensions. Auerbach, Mazur, and Ulam established the conjecture for the case $n=2$ in [1]. Dvoretzky [5] later proved it for infinite-dimensional spaces $V$ and all $n\geq 2$. Gromov [7] showed that the answer is positive whenever $n$ is even or $\dim(V)\geq n+2$. Bor, Hernández-Lamoneda, Jiménez-Desantiago, and Montejano extended Gromov’s theorem to all $n$ congruent to $1$ modulo $4$, with the exception of $n=133$; see [3]. More recently, Ivanov, Mamaev, and Nordskova [8] obtained an affirmative solution for $n=3$. The problem can also be formulated for complex normed spaces; see [4], [7], [10] for further details.
Dvoretzky and Gromov reduced the problem to the situation where $\mbox{dim}(V)=n+1$ for some finite integer $n\geq 2$. Then, the problem can be restated in geometric terms as follows.
Problem 1.2.
Let $n\geq 3$ and let $K\subseteq\mathbb{R}^{n}$ be an origin-symmetric convex body. Assume that all intersections of $K$ with (n1)(n-1)-dimensional linear subspaces are linearly equivalent. Is $K$ neccessarily an ellipsoid?
The symmetry condition follows directly from Montejano in [11]: If $K\subseteq\mathbb{R}^{n}$ is a convex body and all intersections of $K$ with $n$-dimensional linear subspaces are affine equivalent, then either $K$ is symmetric with respect to 0, or $K$ is a (not necessarily centered) ellipsoid. Our principal result is the following theorem, which provides affirmative answers for every finite $n\geq 3$.
Theorem 1.1.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$, if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then $K$ is an ellipsoid.
The proof of Theorem 1.1 is divided into two parts. Assume first that $K\cap\xi_{0}$ is an (n1)(n-1)-dimensional ellipsoid. Under this assumption, $K$ itself is an ellipsoid. Thus, to prove the main theorem, we choose a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then employ the linear invariance of the John ellipsoid to derive the final result.
2. Notation and Preliminaries
This section mainly presents some basic content, such as symbols and definitions. For more detailed content, readers can refer to the books by Gardner [6] and Schneider [12].
Let $\mathbb{R}^{n}$ be the classical $n$-dimensional Euclidean space. We denote by $\|x\|$ the Euclidean norm of $x\in\mathbb{R}^{n}$. We write $B^{n}=\{x\in\mathbb{R}^{n}:\|x\|\leq 1\}$ for the unit ball and $S^{n-1}=\{x\in\mathbb{R}^{n}:\|x\|=1\}$ for the unit sphere. We denote $\{e_{1},e_{2},\dots,e_{n}\}$ to be the standard basis in $\mathbb{R}^{n}$. For any $\xi\in S^{n-1}$, we denote $\xi^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle=0\}$ to be a hyperplane that is perpendicular to $\xi$. Here, $\langle x,\xi\rangle$ is the usual inner product in $\mathbb{R}^{n}$. The Grassmann manifold of $k$-dimensional subspaces in $\mathbb{R}^{n}$ is denoted by $G(n,k)$. $GL(n)$ and $SL(n)$ are the general linear group and the special linear group in $\mathbb{R}^{n}$. For $\phi\in GL(n)$, we denote $|\phi|=|\det(\phi)|$ by the norm of the operator $\phi$.
A set in $\mathbb{R}^{n}$ is called convex if it contains the closed line segment joining any two of its points. A convex set is a convex body if it is compact and has a non-empty interior. A convex body is strictly convex if its boundary contains no line segments.
A compact set $L$ is called a star body if the origin $O$ is an interior point of $L$, every line through $O$ meets $L$ in a line segment, and its Minkowski gauge defined by
$\|x\|_{L}=\inf\{a\geq 0:x\in aL\}$
is a continuous function on $\mathbb{R}^{n}$.
[…]
For $H\in G(n,k)$, we call a body a revolution body with respect to $H$ if, for any special orthonormal transformation $\phi\in SO(n)$ with $\phi(H)=H$, we have $\phi(K)=K$.
Let $K\subseteq\mathbb{R}^{n}$ be a convex body. The John ellipsoid of $K$ is the ellipsoid of maximal volume contained in $K$, denoted by $J(K)$. It is clear that the John ellipsoid is invariant under linear transformations; that is, for any $\phi\in GL(n)$, $J(\phi(K))=\phi(J(K))$. When the John ellipsoid is the $n$-dimensional unit Euclidean ball $B^{n}$, the body $K$ is said to be in John position.
3. Proof of the main theorem
We begin the proof with a special case.
Theorem 3.1.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric star body. If for any $\xi\in\mathbb{S}^{n-1}$, $K\cap\xi^{\perp}$ is a centered ellipsoid, then $K$ is an ellipsoid.
Lemma 3.1.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. If there exists a $H\in G(n,k)$ for $1\leq k\leq n-1$ such that $K\cap H$ and $K\cap H^{\perp}$ are balls, then $K$ is an ellipsoid.
Proof.
Suppose that $K\cap H$ is a $k$-dimensional ball with a radius $r_{1}$, and $K\cap H^{\perp}$ is a (nk)(n-k)-dimensional ball with a radius $r_{2}$.
Then, for any $\eta\in\mathbb{S}^{n-1}\cap H$, $K\cap\eta^{\perp}$ is an ellipsoid where $K\cap\eta^{\perp}\cap H$ is a (k1)(k-1)-dimensional ball with a radius $r_{1}$, and $K\cap\eta^{\perp}\cap H^{\perp}$ is a (nk)(n-k)-dimensional ball with a radius $r_{2}$.
Therefore, for any $\eta\in\mathbb{S}^{n-1}\cap H$, $K\cap\eta^{\perp}$ is an ellipsoid with (k1)(k-1) axes with lengths $r_{1}$ and (nk)(n-k) axes with lengths $r_{2}$, which implies $K$ is a revolution body with respect to $H^{\perp}$; hence, $K$ is an ellipsoid. ∎
Lemma 3.2.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. If we define the maximal set $M(K)$ to be
$M(K)=\{\theta\in\mathbb{S}^{n-1}:\rho_{K}(\theta)\geq\rho_{K}(\eta)\mbox{ for any }\eta\in\mathbb{S}^{n-1}\},$
then, for any linearly independent vectors $\theta_{1},\cdots,\theta_{k}\in M(K)$ with $1\leq k\leq n$, $K\cap\mbox{span}(\theta_{1},\cdots,\theta_{k})$ is a $k$-dimensional ball with a maximal radius.
Proof.
Case I. If $k=1$, $K\cap\mbox{span}(\theta_{1})$ is a $1$-dimensional ball, as $K$ is origin symmetric.
Case II. If $2\leq k\leq n-1$, $K\cap\mbox{span}(\theta_{1},\cdots,\theta_{k})$ is a $k$-dimensional ellipsoid with $k$ independent vectors $\theta_{1},\cdots,\theta_{k}$ where $\rho_{K}(\theta_{1})=\cdots=\rho_{K}(\theta_{k})$.
Hence, $K\cap\mbox{span}(\theta_{1},\cdots,\theta_{k})$ has to be a $k$-dimensional ball with a maximal radius.
Case III. If $k=n$, by Case II, for any $1\leq i\leq n$, $K\cap\,\mbox{span}\left(\{\theta_{j}\}_{j\neq i}\right)$ is a (n1)(n-1)-dimensional ball. For any $\xi\in\mathbb{S}^{n-1}$, if $\xi^{\perp}$ coincides with $\mbox{span}\left(\{\theta_{j}\}_{j\neq i}\right)$ for some $i$, $K\cap\xi^{\perp}$ is a (n1)(n-1)-dimensional ball with a maximal radius. If $\xi^{\perp}$ does not coincide with $\mbox{span}\left(\{\theta_{j}\}_{j\neq i}\right)$ for all $i$, we can choose n1n-1 linearly independent vectors in $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{1}}\right)$ and $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{2}}\right)$ with $i_{1}\neq i_{2}$; otherwise, $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{1}}\right)$ and $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{2}}\right)$ coincide, which gives $i_{1}=i_{2}$.
Therefore, by Case II, $K\cap\xi^{\perp}$ has to be a $k$-dimensional ball with a maximal radius. Therefore, for any $\xi\in\mathbb{S}^{n-1}$, $K\cap\xi^{\perp}$ has to be a $k$-dimensional ball with a maximal radius, which implies that $K$ is a ball. ∎
It is clear to derive the following two Corollaries from Lemma 3.2.
Corollary 3.1.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. Then the maximal set $M(k)$ is a $k$-dimensional sphere in $\mathbb{S}^{n-1}$ for some $0\leq k\leq n-1$.
Corollary 3.2.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. If we define the minimal set $m(K)$ to be
$m(K)=\{\theta\in\mathbb{S}^{n-1}:\rho_{K}(\theta)\leq\rho_{K}(\eta)\mbox{ for any }\eta\in\mathbb{S}^{n-1}\},$
then, $m(k)$ is a $l$-dimensional sphere in $\mathbb{S}^{n-1}$ for some $0\leq l\leq n-1$.
Lemma 3.3.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. Then either the maximal set $M(k)$ is orthogonal to the minimal set $m(K)$ or $M(K)=m(K)=\mathbb{S}^{n-1}$.
Proof.
If $M(K)\neq m(K)$, then $M(K)\cap m(K)=\emptyset$, and thus
$\mbox{dim}(M(K))+\mbox{dim}(m(K))\leq n-2;$
otherwise, $\mbox{dim}(\mbox{span}(M(K)))+\mbox{dim}(\mbox{span}(m(K)))\geq n+1$, where $\mathbb{S}^{n-1}\cap\mbox{span}(M(K))\cap\mbox{span}(m(K))\neq\emptyset$.
Case I. If $M(K)$ is 0-dimensional and $m(K)$ is (n2)(n-2)-dimensional, we choose $\xi\in M(K)$ and consider any (n3)(n-3)-dimensional subsphere $H$ of $m(K)$.
Then $K\cap\mbox{span}(\xi,H)$ forms an ellipsoid, where the ellipsoid attains its maximum at $\xi$ and its minimum in $H$; hence $\xi\perp H$, which gives $M(K)\perp m(K)$.
Case II. If $M(K)$ is $k$-dimensional ($k\geq 1$) and $m(K)$ is $l$-dimensional ($l\leq n-3$), then for any $\xi\in M(K)$, $K\cap\,\mbox{span}(\xi,m(K))$ is an $l+2$-dimensional ellipsoid, where the ellipsoid attains its maximum at $\xi$ and its minimum in $m(K)$; hence $\xi\perp m(K)$, which gives $M(K)\perp m(K)$. ∎
Proof of Theorem 3.1.
First by Corollary 3.1 and 3.2, we can set the maximal set $M(K)$ to be $k_{1}$-dimensional and $m(K)$ to be $l_{1}$-dimensional. If $M(K)=m(K)=\mathbb{S}^{n-1}$, $K$ is a ball. If not, by Lemma 3.3, we have $M(K)\perp m(K)$ with $k_{1}+l_{1}\leq n-2$.
If $k_{1}+l_{1}=n-2$, by Lemma 3.1, we have $K$ to be ellipsoid.
If $k_{1}+l_{1}<n-2$, we can define an orthonomal bases $\{e_{1},\dots,e_{n}\}$ such that $\mathbb{S}^{n-1}\cap\mbox{span}\left(e_{1},\dots,e_{k_{1}+1}\right)=M(K)$ and $\mathbb{S}^{n-1}\cap\mbox{span}\left(e_{n-l_{1}},\dots,e_{n}\right)=m(K)$.
Now we can set $M_{1}:=\max\{\rho_{K}(\theta):\theta\in\mathbb{S}^{n-1}\}$, $m_{1}=\min\{\rho_{K}(\theta):\theta\in\mathbb{S}^{n-1}\}$, and a linear transformation $T_{1}$ to be
$\left(\begin{array}[]{ccc}\frac{m_{1}}{M_{1}}I_{k_{1}+1}&&\\ &I_{n-k_{1}-l_{1}-2}&\\ &&I_{l_{1}+1}.\end{array}\right)$
Take $K_{2}=T_{1}(K)$. It is clear that $K_{2}$ satisfies the given condition in Theorem 3.1, as $T_{1}$ is a linear transformation. Moreover,
$K_{2}\cap\mbox{span}\{e_{1},\dots,e_{k_{1}+1},e_{n-l_{1}},\dots,e_{n}\}$
is a $(k_{1}+l_{1}+2)$-dimensional ball with radius $m_{1}$.
Note that, by Corollary 3.1 and 3.2, we can set the maximal set $M(K_{2})$ to be $k_{2}$-dimensional and $m(K_{2})$ to be $l_{2}$-dimensional.
We claim $\mathbb{S}^{n-1}\cap\mbox{span}(M(K),m(K))\subseteq m(K_{2})$. Hence,
$l_{2}\geq k_{1}+l_{1}+1.$
Then, by Lemma 3.3, either $m(K_{2})=\mathbb{S}^{n-1}$ or $M(K_{2})\perp m(K_{2})$.
If $m(K_{2})=\mathbb{S}^{n-1}$, we have $K_{2}$ to be a ball which gives $K$ being an ellipsoid.
If $M(K_{2})\perp m(K_{2})$, we choose the same orthonomal bases $\{e_{1},\dots,e_{n}\}$ such that $\mathbb{S}^{n-1}\cap\mbox{span}\left(e_{k_{1}+1},\dots,e_{k_{1}+k_{2}+2}\right)=M(K_{2})$.
Now we can set $M_{2}:=\max\{\rho_{K_{2}}(\theta):\theta\in\mathbb{S}^{n-1}\}$, $m_{2}=\min\{\rho_{K_{2}}(\theta):\theta\in\mathbb{S}^{n-1}\}$ with $m_{2}=m_{1}$, and a linear transformation $T_{2}$ to be
$\left(\begin{array}[]{cccc}I_{k_{1}+1}&&&\\ &\frac{m_{1}}{M_{2}}I_{k_{2}+1}&&\\ &&I_{n-k_{1}-l_{1}-k_{2}-3}&\\ &&&I_{l_{1}+1}.\end{array}\right)$
Take $K_{3}=T_{2}(K_{2})$. It is clear that $K_{3}$ satisfies the given condition in Theorem 3.1, as $T_{2}$ is a linear transformation. Moreover,
$K_{3}\cap\mbox{span}\{e_{1},\dots,e_{k_{1}+k_{2}+2},e_{n-l_{1}},\dots,e_{n}\}$
is a $(k_{1}+k_{2}+l_{1}+3)$-dimensional ball with radius $m_{1}$.
Repeat the above process, we can finally get a ball after finite linear transformations, which means $K$ is an ellipsoid.
proof of the claim. Since $m(K_{2})\leq m_{1}$, if we assume the opposite of the claim, $m(K_{2})<m_{1}$, which means there exists $\eta\in\mathbb{S}^{n-1}$ such that $\rho_{K_{2}}(\eta)<m_{1}$ and $\eta\notin M(K)\cap m(K)$.
Thus, we choose the preimage of $\eta$ under $T_{1}$ to be $\eta$’. Since $k_{1}+l_{1}<n-2$, we have $K\cap\mbox{span}(M(K),\eta^{\prime})$ is at most $k_{1}+2$ dimensional with $k_{1}+2\leq n-1$ thus an ellipsoid with longest axes in $M(K)$ and another axis $b$. Note that $m_{1}$ is the minimum of $K$, then $b\geq m(K)$.
Now we decompose $\rho_{K}(\eta^{\prime})\eta^{\prime}$ into two parts
$\rho_{K}(\eta^{\prime})\eta^{\prime}=M_{1}\eta_{1}+\tilde{b}\eta_{2},$
where $\eta_{1}\in M(K)$ and $\eta_{2}\in\mathbb{S}^{n-1}\cap M(K)^{\perp}$.
And thus, we consider the elliptic $K\cap\mbox{span}(\eta_{1},\eta_{2})$ with long axis $M_{1}$ and short axis $b$. Then $T_{1}(K\cap\mbox{span}(\eta_{1},\eta_{2}))$ gives another elliptic with long axis $b\,(\geq m_{1})$ and short axis $m_{1}$, where $\rho_{K_{2}}(\eta)\eta\in T_{1}(K\cap\mbox{span}(\eta_{1},\eta_{2}))$ with $\rho_{K_{2}}(\eta)\geq m_{1}$, contradiction.
Now we can prove the main theorem using John ellipsoids.
Theorem 3.2.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}.$
Moreover, $E=J(K)$.
Proof.
First, we need to show the existence of the star body $E$. It is sufficient to show the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$. Since $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, for any $x\in K\cap\xi_{0}^{\perp}$, we have
$\|\phi_{\xi}(x)\|_{K}$ $=\inf\{\lambda:\phi_{\xi}(x)\in\lambda K\}$
$=\inf\{\lambda:\phi_{\xi}(x)\in\lambda(K\cap\xi^{\perp})\}$
$=\inf\{\lambda:\phi_{\xi}(x)\in\lambda(\phi_{\xi}(K\cap\xi_{0}^{\perp}))\}$
$=\inf\left\{\lambda:\phi_{\xi}(x)\in\phi_{\xi}\left(\frac{\lambda}{|\phi_{\xi}|}(K\cap\xi_{0}^{\perp})\right)\right\}$
$=\inf\left\{\lambda:x\in\frac{\lambda}{|\phi_{\xi}|}(K\cap\xi_{0}^{\perp})\right\}$
$=|\phi_{\xi}|\inf\left\{\lambda:x\in\lambda(K\cap\xi_{0}^{\perp})\right\}$
$=|\phi_{\xi}|\|x\|_{K}.$
Then the continuity of the Minkowski gauge holds
$\left|\|\phi_{\xi}(x)\|_{K}-\|\phi_{\eta}(x)\|_{K}\right|\leq\|x\|_{K}\big||\phi_{\xi}|-|\phi_{\eta}|\big|\leq c\|x\|_{K}\|\xi-\eta\|$
for some constant $c$. The second inequality follows from the continuity of the volume of the sections.
Therefore, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$ follows, which gives an origin-symmetric star body
$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}.$
Then, by Theorem 3.1, $E$ is an ellipsoid in $K$ with $E\cap\xi^{\perp}$ being the John ellipsoid of $K\cap\xi^{\perp}$ for any $\xi\in\mathbb{S}^{n-1}$.
For any ellipsoid $\tilde{E}$ in $K$ and $\xi\in\mathbb{S}^{n-1}$, $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$ because $E\cap\xi^{\perp}$ is the John ellipsoid of $K\cap\xi^{\perp}$. As a result of ellipsoid being the intersection body, by Koldobsky and Zhang’s result (see [9] and [13]), we have $|\tilde{E}|\leq|E|$, which implies $E$ is the John ellipsoid of $K$. ∎
Proof of the Theorem 1.1.
By Theorem 2.2, $K$ contain a John ellipsoid $E$ where for any $\xi\in\mathbb{S}^{n-1}$, $E\cap\xi^{\perp}$ is the John ellipsoid of $K\cap\xi^{\perp}$. We can affine transform $K$ to the John’s position $\tilde{K}$, where $E$ becomes a ball $B$. Since John ellipsoid is linear invariant, $\tilde{K}\cap\xi^{\perp}$ has $B\cap\xi^{\perp}$ to be its John ellipsoid, then $\phi_{\xi}$ has to be congruent transformation; otherwise, the John ellipsoid of $\phi_{\xi}(\tilde{K}\cap\xi_{0}^{\perp})$ is $\phi_{\xi}(B\cap\xi^{\perp})$ which is not $B\cap\xi^{\perp}$.
Then $\tilde{K}\cap\xi^{\perp}$ have some volumes for all $\xi\in\mathbb{S}^{n-1}$, and thus by Minkowski-Funk theorem, $\tilde{K}$ has to be a ball, which implies $K$ is an ellipsoid. ∎

### v2

@@
 In this paper, we first prove that the origin-symmetric star body with ellipsoidal sections is an ellipsoid.
-Then we give a complete proof of Banach’s isometric subspace problem in finite dimensions through the John ellipsoid.
+Then, we provide a complete proof of Banach’s isometric subspace problem in finite dimensions through the John ellipsoid.
+Key words and phrases:
+Ellipsoid characterization, convex body, central sections, John ellipsoid.
 2010 Mathematics Subject Classification:
@@
 Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric star body. If for any $\xi\in\mathbb{S}^{n-1}$, $K\cap\xi^{\perp}$ is a centered ellipsoid, then $K$ is an ellipsoid.
-Lemma 3.1.
-Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1.
-If there exists a $H\in G(n,k)$ for $1\leq k\leq n-1$ such that $K\cap H$ and $K\cap H^{\perp}$ are balls, then $K$ is an ellipsoid.
-Proof.
-Suppose that $K\cap H$ is a $k$-dimensional ball with a radius $r_{1}$, and $K\cap H^{\perp}$ is a (nk)(n-k)-dimensional ball with a radius $r_{2}$.
-Then, for any $\eta\in\mathbb{S}^{n-1}\cap H$, $K\cap\eta^{\perp}$ is an ellipsoid where $K\cap\eta^{\perp}\cap H$ is a (k1)(k-1)-dimensional ball with a radius $r_{1}$, and $K\cap\eta^{\perp}\cap H^{\perp}$ is a (nk)(n-k)-dimensional ball with a radius $r_{2}$.
-Therefore, for any $\eta\in\mathbb{S}^{n-1}\cap H$, $K\cap\eta^{\perp}$ is an ellipsoid with (k1)(k-1) axes with lengths $r_{1}$ and (nk)(n-k) axes with lengths $r_{2}$, which implies $K$ is a revolution body with respect to $H^{\perp}$; hence, $K$ is an ellipsoid.
-∎
-Lemma 3.2.
-Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. If we define the maximal set $M(K)$ to be
-$M(K)=\{\theta\in\mathbb{S}^{n-1}:\rho_{K}(\theta)\geq\rho_{K}(\eta)\mbox{ for any }\eta\in\mathbb{S}^{n-1}\},$
-then, for any linearly independent vectors $\theta_{1},\cdots,\theta_{k}\in M(K)$ with $1\leq k\leq n$, $K\cap\mbox{span}(\theta_{1},\cdots,\theta_{k})$ is a $k$-dimensional ball with a maximal radius.
-Proof.
-Case I. If $k=1$, $K\cap\mbox{span}(\theta_{1})$ is a $1$-dimensional ball, as $K$ is origin symmetric.
-Case II.
-If $2\leq k\leq n-1$, $K\cap\mbox{span}(\theta_{1},\cdots,\theta_{k})$ is a $k$-dimensional ellipsoid with $k$ independent vectors $\theta_{1},\cdots,\theta_{k}$ where $\rho_{K}(\theta_{1})=\cdots=\rho_{K}(\theta_{k})$.
-Hence, $K\cap\mbox{span}(\theta_{1},\cdots,\theta_{k})$ has to be a $k$-dimensional ball with a maximal radius.
-Case III.
-If $k=n$, by Case II, for any $1\leq i\leq n$, $K\cap\,\mbox{span}\left(\{\theta_{j}\}_{j\neq i}\right)$ is a (n1)(n-1)-dimensional ball.
-For any $\xi\in\mathbb{S}^{n-1}$, if $\xi^{\perp}$ coincides with $\mbox{span}\left(\{\theta_{j}\}_{j\neq i}\right)$ for some $i$, $K\cap\xi^{\perp}$ is a (n1)(n-1)-dimensional ball with a maximal radius.
-If $\xi^{\perp}$ does not coincide with $\mbox{span}\left(\{\theta_{j}\}_{j\neq i}\right)$ for all $i$, we can choose n1n-1 linearly independent vectors in $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{1}}\right)$ and $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{2}}\right)$ with $i_{1}\neq i_{2}$; otherwise, $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{1}}\right)$ and $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{2}}\right)$ coincide, which gives $i_{1}=i_{2}$.
-Therefore, by Case II, $K\cap\xi^{\perp}$ has to be a $k$-dimensional ball with a maximal radius.
-Therefore, for any $\xi\in\mathbb{S}^{n-1}$, $K\cap\xi^{\perp}$ has to be a $k$-dimensional ball with a maximal radius, which implies that $K$ is a ball.
-∎
-It is clear to derive the following two Corollaries from Lemma 3.2.
-Corollary 3.1.
-Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. Then the maximal set $M(k)$ is a $k$-dimensional sphere in $\mathbb{S}^{n-1}$ for some $0\leq k\leq n-1$.
-Corollary 3.2.
-Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. If we define the minimal set $m(K)$ to be
-$m(K)=\{\theta\in\mathbb{S}^{n-1}:\rho_{K}(\theta)\leq\rho_{K}(\eta)\mbox{ for any }\eta\in\mathbb{S}^{n-1}\},$
-then, $m(k)$ is a $l$-dimensional sphere in $\mathbb{S}^{n-1}$ for some $0\leq l\leq n-1$.
-Lemma 3.3.
-Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. Then either the maximal set $M(k)$ is orthogonal to the minimal set $m(K)$ or $M(K)=m(K)=\mathbb{S}^{n-1}$.
-Proof.
-If $M(K)\neq m(K)$, then $M(K)\cap m(K)=\emptyset$, and thus
-$\mbox{dim}(M(K))+\mbox{dim}(m(K))\leq n-2;$
-otherwise, $\mbox{dim}(\mbox{span}(M(K)))+\mbox{dim}(\mbox{span}(m(K)))\geq n+1$, where $\mathbb{S}^{n-1}\cap\mbox{span}(M(K))\cap\mbox{span}(m(K))\neq\emptyset$.
-Case I. If $M(K)$ is 0-dimensional and $m(K)$ is (n2)(n-2)-dimensional, we choose $\xi\in M(K)$ and consider any (n3)(n-3)-dimensional subsphere $H$ of $m(K)$.
-Then $K\cap\mbox{span}(\xi,H)$ forms an ellipsoid, where the ellipsoid attains its maximum at $\xi$ and its minimum in $H$; hence $\xi\perp H$, which gives $M(K)\perp m(K)$.
-Case II.
-If $M(K)$ is $k$-dimensional ($k\geq 1$) and $m(K)$ is $l$-dimensional ($l\leq n-3$), then for any $\xi\in M(K)$, $K\cap\,\mbox{span}(\xi,m(K))$ is an $l+2$-dimensional ellipsoid, where the ellipsoid attains its maximum at $\xi$ and its minimum in $m(K)$; hence $\xi\perp m(K)$, which gives $M(K)\perp m(K)$.
-∎
-Proof of Theorem 3.1.
-First by Corollary 3.1 and 3.2, we can set the maximal set $M(K)$ to be $k_{1}$-dimensional and $m(K)$ to be $l_{1}$-dimensional.
-If $M(K)=m(K)=\mathbb{S}^{n-1}$, $K$ is a ball.
-If not, by Lemma 3.3, we have $M(K)\perp m(K)$ with $k_{1}+l_{1}\leq n-2$.
-If $k_{1}+l_{1}=n-2$, by Lemma 3.1, we have $K$ to be ellipsoid.
-If $k_{1}+l_{1}<n-2$, we can define an orthonomal bases $\{e_{1},\dots,e_{n}\}$ such that $\mathbb{S}^{n-1}\cap\mbox{span}\left(e_{1},\dots,e_{k_{1}+1}\right)=M(K)$ and $\mathbb{S}^{n-1}\cap\mbox{span}\left(e_{n-l_{1}},\dots,e_{n}\right)=m(K)$.
-Now we can set $M_{1}:=\max\{\rho_{K}(\theta):\theta\in\mathbb{S}^{n-1}\}$, $m_{1}=\min\{\rho_{K}(\theta):\theta\in\mathbb{S}^{n-1}\}$, and a linear transformation $T_{1}$ to be
-$\left(\begin{array}[]{ccc}\frac{m_{1}}{M_{1}}I_{k_{1}+1}&&\\ &I_{n-k_{1}-l_{1}-2}&\\ &&I_{l_{1}+1}.\end{array}\right)$
-Take $K_{2}=T_{1}(K)$. It is clear that $K_{2}$ satisfies the given condition in Theorem 3.1, as $T_{1}$ is a linear transformation. Moreover,
-$K_{2}\cap\mbox{span}\{e_{1},\dots,e_{k_{1}+1},e_{n-l_{1}},\dots,e_{n}\}$
-is a $(k_{1}+l_{1}+2)$-dimensional ball with radius $m_{1}$.
-Note that, by Corollary 3.1 and 3.2, we can set the maximal set $M(K_{2})$ to be $k_{2}$-dimensional and $m(K_{2})$ to be $l_{2}$-dimensional.
-We claim $\mathbb{S}^{n-1}\cap\mbox{span}(M(K),m(K))\subseteq m(K_{2})$. Hence,
-$l_{2}\geq k_{1}+l_{1}+1.$
-Then, by Lemma 3.3, either $m(K_{2})=\mathbb{S}^{n-1}$ or $M(K_{2})\perp m(K_{2})$.
-If $m(K_{2})=\mathbb{S}^{n-1}$, we have $K_{2}$ to be a ball which gives $K$ being an ellipsoid.
-If $M(K_{2})\perp m(K_{2})$, we choose the same orthonomal bases $\{e_{1},\dots,e_{n}\}$ such that $\mathbb{S}^{n-1}\cap\mbox{span}\left(e_{k_{1}+1},\dots,e_{k_{1}+k_{2}+2}\right)=M(K_{2})$.
-Now we can set $M_{2}:=\max\{\rho_{K_{2}}(\theta):\theta\in\mathbb{S}^{n-1}\}$, $m_{2}=\min\{\rho_{K_{2}}(\theta):\theta\in\mathbb{S}^{n-1}\}$ with $m_{2}=m_{1}$, and a linear transformation $T_{2}$ to be
-$\left(\begin{array}[]{cccc}I_{k_{1}+1}&&&\\ &\frac{m_{1}}{M_{2}}I_{k_{2}+1}&&\\ &&I_{n-k_{1}-l_{1}-k_{2}-3}&\\ &&&I_{l_{1}+1}.\end{array}\right)$
-Take $K_{3}=T_{2}(K_{2})$. It is clear that $K_{3}$ satisfies the given condition in Theorem 3.1, as $T_{2}$ is a linear transformation. Moreover,
-$K_{3}\cap\mbox{span}\{e_{1},\dots,e_{k_{1}+k_{2}+2},e_{n-l_{1}},\dots,e_{n}\}$
-is a $(k_{1}+k_{2}+l_{1}+3)$-dimensional ball with radius $m_{1}$.
-Repeat the above process, we can finally get a ball after finite linear transformations, which means $K$ is an ellipsoid.
-proof of the claim.
-Since $m(K_{2})\leq m_{1}$, if we assume the opposite of the claim, $m(K_{2})<m_{1}$, which means there exists $\eta\in\mathbb{S}^{n-1}$ such that $\rho_{K_{2}}(\eta)<m_{1}$ and $\eta\notin M(K)\cap m(K)$.
-Thus, we choose the preimage of $\eta$ under $T_{1}$ to be $\eta$’.
-Since $k_{1}+l_{1}<n-2$, we have $K\cap\mbox{span}(M(K),\eta^{\prime})$ is at most $k_{1}+2$ dimensional with $k_{1}+2\leq n-1$ thus an ellipsoid with longest axes in $M(K)$ and another axis $b$.
-Note that $m_{1}$ is the minimum of $K$, then $b\geq m(K)$.
-Now we decompose $\rho_{K}(\eta^{\prime})\eta^{\prime}$ into two parts
-$\rho_{K}(\eta^{\prime})\eta^{\prime}=M_{1}\eta_{1}+\tilde{b}\eta_{2},$
-where $\eta_{1}\in M(K)$ and $\eta_{2}\in\mathbb{S}^{n-1}\cap M(K)^{\perp}$.
-And thus, we consider the elliptic $K\cap\mbox{span}(\eta_{1},\eta_{2})$ with long axis $M_{1}$ and short axis $b$.
-Then $T_{1}(K\cap\mbox{span}(\eta_{1},\eta_{2}))$ gives another elliptic with long axis $b\,(\geq m_{1})$ and short axis $m_{1}$, where $\rho_{K_{2}}(\eta)\eta\in T_{1}(K\cap\mbox{span}(\eta_{1},\eta_{2}))$ with $\rho_{K_{2}}(\eta)\geq m_{1}$, contradiction.
+Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [13, Theorem 3.1]).
+The convexity of the body $K$ results from the fact that if we have two points in the body, we can examine the plane they span (with the origin).
+Since the planar section is an ellipsoid by assumption, the entire line segment between $x$ and $y$ is contained in $K$.
 Now we can prove the main theorem using John ellipsoids.
@@
 Proof.
-First, we need to show the existence of the star body $E$.
-It is sufficient to show the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$.
-Since $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, for any $x\in K\cap\xi_{0}^{\perp}$, we have
-$\|\phi_{\xi}(x)\|_{K}$ $=\inf\{\lambda:\phi_{\xi}(x)\in\lambda K\}$
-$=\inf\{\lambda:\phi_{\xi}(x)\in\lambda(K\cap\xi^{\perp})\}$
-$=\inf\{\lambda:\phi_{\xi}(x)\in\lambda(\phi_{\xi}(K\cap\xi_{0}^{\perp}))\}$
-$=\inf\left\{\lambda:\phi_{\xi}(x)\in\phi_{\xi}\left(\frac{\lambda}{|\phi_{\xi}|}(K\cap\xi_{0}^{\perp})\right)\right\}$
-$=\inf\left\{\lambda:x\in\frac{\lambda}{|\phi_{\xi}|}(K\cap\xi_{0}^{\perp})\right\}$
-$=|\phi_{\xi}|\inf\left\{\lambda:x\in\lambda(K\cap\xi_{0}^{\perp})\right\}$
-$=|\phi_{\xi}|\|x\|_{K}.$
-Then the continuity of the Minkowski gauge holds
-$\left|\|\phi_{\xi}(x)\|_{K}-\|\phi_{\eta}(x)\|_{K}\right|\leq\|x\|_{K}\big||\phi_{\xi}|-|\phi_{\eta}|\big|\leq c\|x\|_{K}\|\xi-\eta\|$
-for some constant $c$. The second inequality follows from the continuity of the volume of the sections.
-Therefore, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$ follows, which gives an origin-symmetric star body
-$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}.$
-Then, by Theorem 3.1, $E$ is an ellipsoid in $K$ with $E\cap\xi^{\perp}$ being the John ellipsoid of $K\cap\xi^{\perp}$ for any $\xi\in\mathbb{S}^{n-1}$.
-For any ellipsoid $\tilde{E}$ in $K$ and $\xi\in\mathbb{S}^{n-1}$, $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$ because $E\cap\xi^{\perp}$ is the John ellipsoid of $K\cap\xi^{\perp}$.
-As a result of ellipsoid being the intersection body, by Koldobsky and Zhang’s result (see [9] and [13]), we have $|\tilde{E}|\leq|E|$, which implies $E$ is the John ellipsoid of $K$.
+First, we need to show the existence of the star body $E$. It is sufficient to show the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$.
+Now we define a set-valued function $\lambda:\mathbb{S}^{n-1}\to GL(n)$ to be
+$\lambda(\xi):=\{\phi_{\xi}\in GL(n):K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})\}.$
+And we define the linear symmetry group $G_{K\cap\xi_{0}^{\perp}}$ to be
+$G_{K\cap\xi_{0}^{\perp}}:=\{g\in GL(n):g(K\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\}$
+Note that for any $\phi_{\xi},\psi_{\xi}\in\lambda(\xi)$, we have
+$\psi_{\xi}^{-1}\phi_{\xi}(K\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp},$
+which gives $\phi_{\xi}\in\psi_{\xi}\circ G_{K\cap\xi_{0}^{\perp}}$; that is
+$\lambda(\xi)=\phi_{\xi}\circ G_{K\cap\xi_{0}^{\perp}},$
+for some $\phi_{\xi}\in\lambda(\xi)$.
+And thus, the right action of $G_{K\cap\xi_{0}^{\perp}}$ on $GL(n)$ gives the quotient $GL(n)/G_{K\cap\xi_{0}^{\perp}}$, where the set-valued function $\lambda(x)$ becomes a map $\Phi:\mathbb{S}^{n-1}\to GL(n)\backslash G_{K\cap\xi_{0}^{\perp}}$ with
+$K\cap\xi^{\perp}=\Phi(\xi)(K\cap\xi_{0}^{\perp})$
+Now we need to show $\Phi$ is continuous. Note that the operator norm of $\phi_{\xi}\in\Phi(\xi)$ is bounded, since
+$\|\phi_{\xi}\|\leq\frac{\max_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}{\min_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}.$
+Then we can find an arbitary convergent sequence $\{\xi_{n}\}_{n=1}^{\infty}$ with $\lim_{n\to\infty}\xi_{n}=\xi$ and pick $\phi_{\xi_{n}}\in\Phi(\xi_{n})$ such that there exists a subsequence $\{\xi_{n_{k}}\}_{k=1}^{\infty}$ with $\phi_{\xi_{n_{k}}}$ converging to $\phi\in GL(n)$ in operator norm.
+And thus, we have
+$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})$
+$\leq$ $d_{H}(\phi(K\cap\xi_{0}),K\cap\xi_{n_{k}}^{\perp})+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp})$
+$\leq$ $d_{H}(\phi(K\cap\xi_{0}),\phi_{\xi_{n_{k}}}(K\cap\xi_{0}^{\perp}))+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp})$
+$\leq$ $c\left|\|\phi\|_{op}-\|\phi_{\xi_{n_{k}}}\|_{op}\right|+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp}).$
+As $k\to\infty$, we have
+$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})=0,$
+which implies $\phi(K\cap\xi_{0})=K\cap\xi^{\perp}$, that is, $\phi\in\Phi(\xi)$. Therefore, $\Phi(\xi_{n_{k}})\to\Phi(\xi)$ as $k\to\infty$.
+This means that for any subsequence $\{\Phi(\xi_{n_{k}})\}_{k=1}^{\infty}$, there exists a sub-subsequence $\{\Phi(\xi_{n_{k_{l}}})\}_{l=1}^{\infty}$ that converges to $\Phi(\xi)$.
+Consequently, we conclude that $\Phi(\xi_{n})\to\Phi(\xi)$.
+Moreover, for any $g\in G_{K\cap\xi_{0}^{\perp}}$, $g(J(K\cap\xi_{0}^{\perp}))=J(g(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, where the John ellipsoid of $K\cap\xi_{0}^{\perp}$ is unique.
+Hence, $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$.
+Therefore, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$ follows from the continuity of $\Phi(\xi)$.
+Now, we need to show that for any $\xi,\eta\in\mathbb{S}^{n-1}$,
+$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
+holds for any $\theta\in\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\eta^{\perp}$.
+It is sufficient to show for any $\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ and any $\zeta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
+$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
+for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$.
+Here we choose $\phi_{\zeta}\in\Phi(\zeta)$ with
+$K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp}.$
+We partition $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ into two subsets, $A_{1}$ and $A_{2}$.
+The set $A_{1}$ consists of all $\xi$ such that, for every $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$,
+$\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}=g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
+for some $g\in G_{K\cap\xi_{0}^{\perp}}$. The set $B$ consists of all $\xi$ for which there exists a vector $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ such that
+$\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}\neq g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
+for every $g\in G_{K\cap\xi_{0}^{\perp}}$. It is clear that $A_{1}\cap A_{2}=\emptyset$. Moreover, the closedness of $A_{1}$ follows the closedness of $G_{K\cap\xi_{0}^{\perp}}$.
+For $\xi\in A$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
+$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
+where
+$J(K\cap\xi_{0}^{\perp})\cap\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=J(K\cap\xi_{0}^{\perp})\cap g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
+which implies
+$J(K\cap\xi_{0}^{\perp})\cap\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(g(J(K\cap\xi_{0}^{\perp}))\cap K\cap\xi^{\perp}\cap\eta^{\perp}),$
+where $g(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, and thus, we have
+$\phi_{\zeta}(J(K\cap\xi_{0}^{\perp})\cap\phi_{\zeta}^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}))=g\circ g^{-1}(J(K\cap\xi_{0}^{\perp})\cap K\cap\xi^{\perp}\cap\eta^{\perp}),$
+that is,
+$\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))\cap\xi_{0}^{\perp}=J(K\cap\xi_{0}^{\perp})\cap\zeta^{\perp}.$
+Therefore, we obtain
+$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
+for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$.
+For $\xi\in A_{2}$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
+$K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp},$
+where
+$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\zeta}(\xi_{0}^{\perp}).$
+Therefore, using the continuity of $\Phi$ and the fact that $\phi^{-1}_{\zeta}\neq g$, for any
+$\eta=\frac{(1-t)\xi_{0}+t\eta_{\zeta}}{\|(1-t)\xi_{0}+t\eta_{\zeta}\|}\quad\text{with }t\in[0,1],\ \eta_{\zeta}\perp\phi^{-1}_{\zeta}(\xi_{0}^{\perp}),\ \text{and }\eta_{\zeta}\perp\xi_{0},$
+there exists a linear map $\phi_{\eta}$ satisfying
+$\phi_{\eta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$
+and $\phi_{\eta}(\xi^{\perp})=\xi_{0}^{\perp}$. Therefore, we obtain
+$\phi^{-1}_{\eta}(K\cap\xi_{0}^{\perp}\cap\eta^{\perp})=K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\eta}(\xi_{0}^{\perp}),$
+and thus, $\eta\in B$.
+Since $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ is compact, we can choose $\zeta_{0}$ such that $\|\eta_{\zeta_{0}},\xi_{0}\|$ reaches the maximum; then $\eta=\frac{(1-t)\xi_{0}+t\eta_{\zeta_{0}}}{\|(1-t)\xi_{0}+t\eta_{\zeta_{0}}\|}$ is in $A_{2}$.
+Therefore, for any open set $O\subseteq A_{2}$ and any $\xi,\eta\in O$, there exists a linear isomorphism between $K\cap\xi_{0}^{\perp}\cap\xi^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$.
+If this were not the case, then by the continuity of $\Phi$ we could find some $\zeta\in O$ such that
+$\phi_{\alpha}\bigl(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}\bigr)=g\bigl(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}\bigr)$
+for every $\alpha\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi_{0},\zeta)$, where $g\in G_{K\cap\xi_{0}^{\perp}}$. Fixing $\xi\in O$, consider the set
+$\{\eta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:K\cap\xi_{0}^{\perp}\cap\eta^{\perp}=\phi_{\eta}(K\cap\xi_{0}^{\perp}\cap\xi^{\perp}),\,\phi\in GL(n)\},$
+which is closed.
+Consequently, for all $\xi,\eta\in\bar{O}$ there is a linear isomorphism between $K\cap\xi_{0}^{\perp}\cap\xi^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$.
+Then $\bar{O}\subseteq A_{2}$, which shows that $A_{2}$ is closed.
+By the connectedness of $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we obtain either $A_{1}$ or $A_{2}$ as $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$.
+If $A_{1}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, the origin-symmetric star body
+$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}$
+exsits. Then, by Theorem 3.1, $E$ is an ellipsoid in $K$ with $E\cap\xi^{\perp}$ being the John ellipsoid of $K\cap\xi^{\perp}$ for any $\xi\in\mathbb{S}^{n-1}$.
+If $A_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, then any (n2)(n-2)-dimensional sections of $K\cap\xi_{0}^{\perp}$ are mutually linearly dependent; then, using the linear dependence of all (n1)(n-1)-dimensional sections, we deduce that all (n2)(n-2)-dimensional sections of $K$ are likewise linearly dependent; therefore, by Gromov’s theorem [7], $K$ must be an ellipsoid.
+And hence, $E=K$ is an ellipsoid.
+For any ellipsoid $\tilde{E}$ contained in $K$ and any $\xi\in\mathbb{S}^{n-1}$, we have $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$, since $E\cap\xi^{\perp}$ serves as the John ellipsoid of $K\cap\xi^{\perp}$.
+As a result of the ellipsoid being an intersection body, Koldobsky and Zhang’s result (see [9] and [14]) yields $|\tilde{E}|\leq|E|$, and hence $E$ must be the John ellipsoid of $K$.
 ∎
+Remark 3.1.
+The proof of the continuity of $\phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ can also be found in [3, Lemma 1.5].
 Proof of the Theorem 1.1.
-By Theorem 2.2, $K$ contain a John ellipsoid $E$ where for any $\xi\in\mathbb{S}^{n-1}$, $E\cap\xi^{\perp}$ is the John ellipsoid of $K\cap\xi^{\perp}$.
-We can affine transform $K$ to the John’s position $\tilde{K}$, where $E$ becomes a ball $B$.
-Since John ellipsoid is linear invariant, $\tilde{K}\cap\xi^{\perp}$ has $B\cap\xi^{\perp}$ to be its John ellipsoid, then $\phi_{\xi}$ has to be congruent transformation; otherwise, the John ellipsoid of $\phi_{\xi}(\tilde{K}\cap\xi_{0}^{\perp})$ is $\phi_{\xi}(B\cap\xi^{\perp})$ which is not $B\cap\xi^{\perp}$.
-Then $\tilde{K}\cap\xi^{\perp}$ have some volumes for all $\xi\in\mathbb{S}^{n-1}$, and thus by Minkowski-Funk theorem, $\tilde{K}$ has to be a ball, which implies $K$ is an ellipsoid. ∎
+By Theorem 3.2, the body $K$ contains a John ellipsoid $E$ such that, for every $\xi\in\mathbb{S}^{n-1}$, the section $E\cap\xi^{\perp}$ is the John ellipsoid of the section $K\cap\xi^{\perp}$.
+By applying an affine transformation, we can move $K$ into John’s position, denoted by $\tilde{K}$, in which $E$ is transformed into a Euclidean ball $B$.
+Since the John ellipsoid is invariant under linear transformation, the set $\tilde{K}\cap\xi^{\perp}$ has $B\cap\xi^{\perp}$ as its John ellipsoid.
+Consequently, $\phi_{\xi}$ must be a congruence transformation; otherwise, the John ellipsoid of $\phi_{\xi}(\tilde{K}\cap\xi_{0}^{\perp})$ would be $\phi_{\xi}(B\cap\xi^{\perp})$, which differs from $B\cap\xi^{\perp}$.
+Hence, $\tilde{K}\cap\xi^{\perp}$ has the same volume for every $\xi\in\mathbb{S}^{n-1}$, and therefore, by the Minkowski–Funk theorem, $\tilde{K}$ must be a ball, which in turn implies that $K$ is an ellipsoid.
+∎
+4. Acknowledgement
+The crucial idea emerged when the author visited the Tianyuan Mathematical Research Center in Kunming.

### v3

@@
 Abstract.
-In this paper, we first prove that the origin-symmetric star body with ellipsoidal sections is an ellipsoid.
-Then, we provide a complete proof of Banach’s isometric subspace problem in finite dimensions through the John ellipsoid.
+In this paper, we begin by constructing an ellipsoid derived from the John ellipsoid of the sections.
+We then use this framework to present a full proof of Banach’s isometric subspace problem in finite-dimensional spaces via the John ellipsoid.
 Key words and phrases:
@@
 More recently, Ivanov, Mamaev, and Nordskova [8] obtained an affirmative solution for $n=3$.
-The problem can also be formulated for complex normed spaces; see [4], [7], [10] for further details.
+The problem can also be formulated for complex normed spaces; see [4], [7], and [10] for further details.
 Dvoretzky and Gromov reduced the problem to the situation where $\mbox{dim}(V)=n+1$ for some finite integer $n\geq 2$. Then, the problem can be restated in geometric terms as follows.
@@
 Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$, if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then $K$ is an ellipsoid.
-The proof of Theorem 1.1 is divided into two parts.
-Assume first that $K\cap\xi_{0}$ is an (n1)(n-1)-dimensional ellipsoid.
-Under this assumption, $K$ itself is an ellipsoid.
-Thus, to prove the main theorem, we choose a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then employ the linear invariance of the John ellipsoid to derive the final result.
+The proof of Theorem 1.1 is divided into two steps.
+We first select a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then use the linear invariance property of John ellipsoids to obtain a corresponding ellipsoid inside $K$.
+With this large ellipsoid at hand, we then establish the desired conclusion.
 2. Notation and Preliminaries
@@
 $GL(n)$ and $SL(n)$ are the general linear group and the special linear group in $\mathbb{R}^{n}$.
-For $\phi\in GL(n)$, we denote $|\phi|=|\det(\phi)|$ by the norm of the operator $\phi$.
+For $\phi\in GL(n)$, we denote $\|\phi\|_{op}=\max_{x\in B^{n}}\|\phi(x)\|$ by the operator norm of $\phi$.
 A set in $\mathbb{R}^{n}$ is called convex if it contains the closed line segment joining any two of its points.
@@
 Since the planar section is an ellipsoid by assumption, the entire line segment between $x$ and $y$ is contained in $K$.
-Now we can prove the main theorem using John ellipsoids.
-Theorem 3.2.
-Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body.
+Now we can prove the main theorem using John ellipsoids. First, we should introduce a couple of Lemmas.
+Lemma 3.1.
+Let $K\subseteq\mathbb{R}^{n}$ and $\phi\in GL(n)$. Then, for every $x\in\mathbb{R}^{n}$ we have
+$\|x\|_{K}=\|\phi(x)\|_{\phi(K)}.$
+Proof.
+$\|\phi(x)\|_{\phi(K)}$ $=\inf\{\lambda:\phi(x)\in\lambda\phi(K)\}$
+$=\inf\{\lambda:\phi(x)\in\phi(\lambda K)\}$
+$=\inf\{\lambda:x\in\lambda K\}$
+$=\|x\|_{K}.$
+Lemma 3.2.
+Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body.
 Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
-$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}.$
-Moreover, $E=J(K)$.
+$E:=\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}(J(K\cap\xi_{0}^{\perp})).$
 Proof.
-First, we need to show the existence of the star body $E$. It is sufficient to show the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$.
+By the definition of $E$, the set $E$ is an origin-symmetric star set whose radial function is given by
+$\rho_{E}(\theta)=\max\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta):\xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\}.$
+Therefore, to prove that $\rho_{E}(\theta)$ is continuous, it suffices to establish the continuity of the mapping $\xi\mapsto\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$.
 Now we define a set-valued function $\lambda:\mathbb{S}^{n-1}\to GL(n)$ to be
@@
 Now we need to show $\Phi$ is continuous. Note that the operator norm of $\phi_{\xi}\in\Phi(\xi)$ is bounded, since
-$\|\phi_{\xi}\|\leq\frac{\max_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}{\min_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}.$
-Then we can find an arbitary convergent sequence $\{\xi_{n}\}_{n=1}^{\infty}$ with $\lim_{n\to\infty}\xi_{n}=\xi$ and pick $\phi_{\xi_{n}}\in\Phi(\xi_{n})$ such that there exists a subsequence $\{\xi_{n_{k}}\}_{k=1}^{\infty}$ with $\phi_{\xi_{n_{k}}}$ converging to $\phi\in GL(n)$ in operator norm.
+$\|\phi_{\xi}\|_{op}\leq\frac{\max_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}{\min_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}.$
+We may choose an arbitrary sequence $\{\xi_{n}\}_{n=1}^{\infty}$ converging to $\xi$, i.e., with $\lim_{n\to\infty}\xi_{n}=\xi$, and for each $n$ select $\phi_{\xi_{n}}\in\Phi(\xi_{n})$.
+Then there exists a subsequence $\{\xi_{n_{k}}\}_{k=1}^{\infty}$ such that the corresponding operators $\phi_{\xi_{n_{k}}}$ converge in operator norm to some $\phi\in GL(n)$.
 And thus, we have
@@
 As $k\to\infty$, we have
-$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})=0,$
-which implies $\phi(K\cap\xi_{0})=K\cap\xi^{\perp}$, that is, $\phi\in\Phi(\xi)$. Therefore, $\Phi(\xi_{n_{k}})\to\Phi(\xi)$ as $k\to\infty$.
+$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})=0.$
+This implies that $\phi(K\cap\xi_{0})=K\cap\xi^{\perp}$, and hence $\phi\in\Phi(\xi)$. Consequently, we have $\Phi(\xi_{n_{k}})\to\Phi(\xi)$ as $k\to\infty$.
 This means that for any subsequence $\{\Phi(\xi_{n_{k}})\}_{k=1}^{\infty}$, there exists a sub-subsequence $\{\Phi(\xi_{n_{k_{l}}})\}_{l=1}^{\infty}$ that converges to $\Phi(\xi)$.
 Consequently, we conclude that $\Phi(\xi_{n})\to\Phi(\xi)$.
-Moreover, for any $g\in G_{K\cap\xi_{0}^{\perp}}$, $g(J(K\cap\xi_{0}^{\perp}))=J(g(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, where the John ellipsoid of $K\cap\xi_{0}^{\perp}$ is unique.
-Hence, $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$.
-Therefore, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$ follows from the continuity of $\Phi(\xi)$.
-Now, we need to show that for any $\xi,\eta\in\mathbb{S}^{n-1}$,
+Moreover, for any $g\in G_{K\cap\xi_{0}^{\perp}}$, we have
+$g(J(K\cap\xi_{0}^{\perp}))=J(g(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp}),$
+since the John ellipsoid of $K\cap\xi_{0}^{\perp}$ is unique.
+Consequently, $J(K\cap\xi_{0}^{\perp})$ is invariant under the group $G_{K\cap\xi_{0}^{\perp}}$.
+Thus, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ as a function of $\xi$ follows from the continuity of $\Phi(\xi)$.
+∎
+Remark 3.1.
+The proof of the continuity of $\phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ can also be found in [3, Lemma 1.5].
+Now we consider the origin-symmetric star body $E$. The next step is to confirm that $E\cap\xi^{\perp}$ indeed recovers $\phi_{\xi}(J(E\cap\xi_{0}^{\perp}))$.
+Lemma 3.3.
+Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body.
+Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then one of the following conditions must be satisfied.
+  1. (1)
+    For all $\xi,\eta\in\mathbb{S}^{n-1}$ we have
     $\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
-holds for any $\theta\in\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\eta^{\perp}$.
-It is sufficient to show for any $\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ and any $\zeta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
+    for every $\theta\in\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\eta^{\perp}$.
+  2. (2)
+    The body $K$ is an ellipsoid.
+Proof.
+Condition (1) can be reformulated as follows: for any $\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ and any $\zeta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we require that
 $\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
-for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$.
-Here we choose $\phi_{\zeta}\in\Phi(\zeta)$ with
+for every $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$, since, by choosing $\eta$ to be a basic direction, we have
+$\phi_{\xi}(K\cap\xi^{\perp})=\phi_{\xi}\phi^{-1}_{\eta}(K\cap\eta^{\perp})$
+for every $\xi\in\mathbb{S}^{n-1}$.
+First, we introduce a function $\Gamma:\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\to\mathbb{R}_{\geq 0}$ defined by
+$\Gamma(\xi)=\sup\Big\{\Big|\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|:\ \zeta,\eta\in\mathbb{S}^{n-1}\cap\,\mathrm{span}(\xi,\xi_{0})\Big\}.$
+Then $\Gamma(\xi)$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$.
+Furthermore, $\Gamma(\xi)=\Gamma(\eta)$ whenever there exists a linear transformation between $K\cap\xi_{0}^{\perp}\cap\xi$ and $K\cap\xi_{0}^{\perp}\cap\eta$, as implied by Lemma 3.1.
+Now fix $\xi$. For any $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$, select $\phi_{\zeta}\in\Phi(\zeta)$ such that
 $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp}.$
@@
 $\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}=g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
-for some $g\in G_{K\cap\xi_{0}^{\perp}}$. The set $B$ consists of all $\xi$ for which there exists a vector $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ such that
+for some $g\in G_{K\cap\xi_{0}^{\perp}}$. The set $A_{2}$ consists of all $\xi$ for which there exists a vector $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ such that
 $\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}\neq g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
-for every $g\in G_{K\cap\xi_{0}^{\perp}}$. It is clear that $A_{1}\cap A_{2}=\emptyset$. Moreover, the closedness of $A_{1}$ follows the closedness of $G_{K\cap\xi_{0}^{\perp}}$.
-For $\xi\in A$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
+for every $g\in G_{K\cap\xi_{0}^{\perp}}$. It is clear that $A_{1}\cap A_{2}=\emptyset$.
+We once more partition $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ into two subsets, $B_{1}$ and $B_{2}$. Define
+$B_{1}:=\{\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:\Gamma(\xi)=0\},\quad B_{2}:=\{\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:\Gamma(\xi)>0\}.$
+Our goal is to show that $B_{1}$ and $B_{2}$ are closed.
+Note that, for $\xi\in A_{1}$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
 $\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
@@
 $\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
-for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$.
-For $\xi\in A_{2}$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
+for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$. And thus, $\Gamma(\xi)=0$, so $A_{1}\subseteq B_{1}$, which implies $B_{2}\subseteq A_{2}$.
+We now focus on the set $B_{2}$.
+For any $\xi\in B_{2}$, and therefore $\xi\in A_{2}$, and for any $\zeta,\eta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$ and $\zeta\neq\eta$, we have
 $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp},$
-where
-$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\zeta}(\xi_{0}^{\perp}).$
-Therefore, using the continuity of $\Phi$ and the fact that $\phi^{-1}_{\zeta}\neq g$, for any
-$\eta=\frac{(1-t)\xi_{0}+t\eta_{\zeta}}{\|(1-t)\xi_{0}+t\eta_{\zeta}\|}\quad\text{with }t\in[0,1],\ \eta_{\zeta}\perp\phi^{-1}_{\zeta}(\xi_{0}^{\perp}),\ \text{and }\eta_{\zeta}\perp\xi_{0},$
-there exists a linear map $\phi_{\eta}$ satisfying
-$\phi_{\eta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$
-and $\phi_{\eta}(\xi^{\perp})=\xi_{0}^{\perp}$. Therefore, we obtain
-$\phi^{-1}_{\eta}(K\cap\xi_{0}^{\perp}\cap\eta^{\perp})=K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\eta}(\xi_{0}^{\perp}),$
-and thus, $\eta\in B$.
-Since $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ is compact, we can choose $\zeta_{0}$ such that $\|\eta_{\zeta_{0}},\xi_{0}\|$ reaches the maximum; then $\eta=\frac{(1-t)\xi_{0}+t\eta_{\zeta_{0}}}{\|(1-t)\xi_{0}+t\eta_{\zeta_{0}}\|}$ is in $A_{2}$.
-Therefore, for any open set $O\subseteq A_{2}$ and any $\xi,\eta\in O$, there exists a linear isomorphism between $K\cap\xi_{0}^{\perp}\cap\xi^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$.
-If this were not the case, then by the continuity of $\Phi$ we could find some $\zeta\in O$ such that
-$\phi_{\alpha}\bigl(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}\bigr)=g\bigl(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}\bigr)$
-for every $\alpha\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi_{0},\zeta)$, where $g\in G_{K\cap\xi_{0}^{\perp}}$. Fixing $\xi\in O$, consider the set
-$\{\eta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:K\cap\xi_{0}^{\perp}\cap\eta^{\perp}=\phi_{\eta}(K\cap\xi_{0}^{\perp}\cap\xi^{\perp}),\,\phi\in GL(n)\},$
-which is closed.
-Consequently, for all $\xi,\eta\in\bar{O}$ there is a linear isomorphism between $K\cap\xi_{0}^{\perp}\cap\xi^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$.
-Then $\bar{O}\subseteq A_{2}$, which shows that $A_{2}$ is closed.
-By the connectedness of $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we obtain either $A_{1}$ or $A_{2}$ as $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$.
-If $A_{1}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, the origin-symmetric star body
-$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}$
-exsits. Then, by Theorem 3.1, $E$ is an ellipsoid in $K$ with $E\cap\xi^{\perp}$ being the John ellipsoid of $K\cap\xi^{\perp}$ for any $\xi\in\mathbb{S}^{n-1}$.
-If $A_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, then any (n2)(n-2)-dimensional sections of $K\cap\xi_{0}^{\perp}$ are mutually linearly dependent; then, using the linear dependence of all (n1)(n-1)-dimensional sections, we deduce that all (n2)(n-2)-dimensional sections of $K$ are likewise linearly dependent; therefore, by Gromov’s theorem [7], $K$ must be an ellipsoid.
-And hence, $E=K$ is an ellipsoid.
-For any ellipsoid $\tilde{E}$ contained in $K$ and any $\xi\in\mathbb{S}^{n-1}$, we have $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$, since $E\cap\xi^{\perp}$ serves as the John ellipsoid of $K\cap\xi^{\perp}$.
-As a result of the ellipsoid being an intersection body, Koldobsky and Zhang’s result (see [9] and [14]) yields $|\tilde{E}|\leq|E|$, and hence $E$ must be the John ellipsoid of $K$.
+and, in addition,
+$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\eta^{\perp})=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\zeta}(\eta^{\perp}).$
+Hence, by exploiting the linearity of $\phi^{-1}_{\zeta}$, for any $y\in\phi^{-1}_{\zeta}(\eta^{\perp})$, we obtain $\langle\phi_{\zeta}(y),\eta\rangle=0$.
+Equivalently, this can be written as $y^{T}A^{T}_{\phi_{\zeta}}\eta=0$, which implies that $y\in(A^{T}_{\phi_{\zeta}}\eta)^{\perp}$, where $A_{\phi_{\zeta}}$ denotes the matrix associated with the linear transformation $\phi_{\zeta}$.
+Therefore, there exists a linear transform between $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap(A^{T}_{\phi_{\zeta}}\eta)^{\perp}$.
+Furthermore, the set
+$E_{\zeta,\xi}:=\{A^{T}_{\phi_{\zeta}}\eta:\eta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})\}$
+forms an ellipse, and the set
+$C_{\zeta,\xi}:=\left\{\frac{A^{T}_{\phi_{\zeta}}\eta}{|A^{T}_{\phi_{\zeta}}\eta|}:\eta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})\right\}$
+gives a circle.
+And thus, there exists a linear transform between $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\tilde{\eta}^{\perp}$ for any $\tilde{\eta}\in C_{\zeta,\xi}$.
+Finally, we have
+[…]
+for any $\tilde{\eta}\in C_{\zeta,\xi}$.
+Note that for any other point $\tilde{\xi}\in B_{2}$, there is a circle $C_{\tilde{\zeta},\tilde{\xi}}$ such that
+$\Gamma(\tilde{\xi})=\Gamma(\bar{\eta})$
+for all $\bar{\eta}\in C_{\tilde{\zeta},\tilde{\xi}}$.
+Consequently, the function $\Gamma$ must be constant on $B_{2}$, since any two circles on $\mathbb{S}^{n-1}$ intersect.
+Therefore, $B_{2}$ is closed.
+By the connectedness of $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we obtain either $B_{1}$ or $B_{2}$ as $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$.
+If $B_{1}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we have $\Gamma(\xi)=0$ for any $\xi$, and hence Condition (1) holds.
+If $B_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ holds, then $A_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ implies that any (n2)(n-2)-dimensional sections of $K\cap\xi_{0}^{\perp}$ are all linearly dependent, because for every $\tilde{\eta}\in C_{\zeta,\xi}$ there is a linear transformation mapping $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ onto $K\cap\xi_{0}^{\perp}\cap\tilde{\eta}^{\perp}$.
+Using the fact that all (n1)(n-1)-dimensional sections are linearly dependent, we then conclude that every (n2)(n-2)-dimensional section of $K$ is also linearly dependent.
+Consequently, by Gromov’s theorem [7], $K$ must be an ellipsoid.
 ∎
-Remark 3.1.
-The proof of the continuity of $\phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ can also be found in [3, Lemma 1.5].
+To establish that $E\cap\xi^{\perp}=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$, it suffices to prove that $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ and $\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))$ coincide for any $k$-dimensional subspace ($1\leq k\leq n-2$).
+This leads to the following lemma, which is proved by induction.
+Lemma 3.4.
+Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body.
+Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then one of the following conditions must be satisfied.
+  1. (1)
+    For all $H\in G(n,k)$ with $1\leq k\leq n-2$, we have
+    $\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
+    for every $\theta\in\mathbb{S}^{n-1}\cap H$ and $\xi,\eta\in H^{\perp}$.
+  2. (2)
+    The body $K$ is an ellipsoid.
+Proof.
+If $E$ is not an ellipsoid, then by Lemma 3.3, Condition (1) is satisfied for k=n2k=n-2. Now assume that Condition (1) holds for $k$.
+For any $H\in G(n,k-1)$ and any $\xi\in H^{\perp}$, we have
+$\mathrm{span}(H,\xi)\in G(n,k).$
+Therefore, by the assumption of Condition (1), we obtain
+$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
+for every $\theta\in\mathbb{S}^{n-1}\cap\,\mathrm{span}(H,\xi)$ and every $\zeta,\eta\in\bigl(\mathrm{span}(H,\xi)\bigr)^{\perp}$. In particular, this implies that
+$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
+for all $\theta\in\mathbb{S}^{n-1}\cap H$ and all $\zeta,\eta\in\bigl(\mathrm{span}(H,\xi)\bigr)^{\perp}$, where
+$\dim\bigl((\mathrm{span}(H,\xi))^{\perp}\bigr)=n-1-\dim(H)\geq 2.$
+That Condition (1) is valid for k1k-1 follows from the fact that
+$(\mathrm{span}(H,\xi_{1}))^{\perp}\cap(\mathrm{span}(H,\xi_{2}))^{\perp}\neq\emptyset$
+for any $\xi_{1},\xi_{2}\in H^{\perp}$.
+In particular, in the case $k=1$, for any $\theta\in\mathbb{S}^{n-1}$ and any $\xi,\eta\in\mathbb{S}^{n-1}\cap\theta^{\perp}$, we have
+$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{E}(\theta).$
+Consequently,
+$E\cap\xi^{\perp}=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi^{\perp}).$
+Now we can show the following theorem.
+Theorem 3.2.
+Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body.
+Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
+$E:=\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}(J(K\cap\xi_{0}^{\perp})).$
+Moreover, $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$ for any $\xi\in\mathbb{S}^{n-1}$ and $E=J(K)$.
+Proof.
+By Lemmas 3.2 and 3.4, either $K$ itself is an ellipsoid with $E=K$, or $E$ is an origin-symmetric star body satisfying $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$.
+In the latter case, Theorem 3.1 implies that $E$ is an ellipsoid.
+Now take an arbitrary ellipsoid $\tilde{E}$ contained in $K$.
+For every $\xi\in\mathbb{S}^{n-1}$, we have the inequality $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$, because $E\cap\xi^{\perp}$ is the John ellipsoid of the section $K\cap\xi^{\perp}$.
+Since the ellipsoid is an intersection body, the theorem of Koldobsky and Zhang (see [9] and [14]) implies that $|\tilde{E}|\leq|E|$.
+Consequently, $E$ is indeed the John ellipsoid of $K$.
+∎
 Proof of the Theorem 1.1.

### v4

@@
 Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$, if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then $K$ is an ellipsoid.
-The proof of Theorem 1.1 is divided into two steps.
-We first select a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then use the linear invariance property of John ellipsoids to obtain a corresponding ellipsoid inside $K$.
+The proof of Theorem 1.1 is divided into three steps.
+We first select a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then use the linear invariance property of John ellipsoids to obtain a corresponding star body $E$ inside $K$.
+Secondly, we use the level set method (see [12, 16, 17] for details) to show $E$ an ellipsoid.
 With this large ellipsoid at hand, we then establish the desired conclusion.
@@
 Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric star body. If for any $\xi\in\mathbb{S}^{n-1}$, $K\cap\xi^{\perp}$ is a centered ellipsoid, then $K$ is an ellipsoid.
-Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [13, Theorem 3.1]).
+Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [14, Theorem 3.1]).
 The convexity of the body $K$ results from the fact that if we have two points in the body, we can examine the plane they span (with the origin).
@@
 Proof.
-Condition (1) can be reformulated as follows: for any $\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ and any $\zeta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we require that
-$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
-for every $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$, since, by choosing $\eta$ to be a basic direction, we have
-$\phi_{\xi}(K\cap\xi^{\perp})=\phi_{\xi}\phi^{-1}_{\eta}(K\cap\eta^{\perp})$
-for every $\xi\in\mathbb{S}^{n-1}$.
-First, we introduce a function $\Gamma:\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\to\mathbb{R}_{\geq 0}$ defined by
-$\Gamma(\xi)=\sup\Big\{\Big|\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|:\ \zeta,\eta\in\mathbb{S}^{n-1}\cap\,\mathrm{span}(\xi,\xi_{0})\Big\}.$
-Then $\Gamma(\xi)$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$.
-Furthermore, $\Gamma(\xi)=\Gamma(\eta)$ whenever there exists a linear transformation between $K\cap\xi_{0}^{\perp}\cap\xi$ and $K\cap\xi_{0}^{\perp}\cap\eta$, as implied by Lemma 3.1.
-Now fix $\xi$. For any $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$, select $\phi_{\zeta}\in\Phi(\zeta)$ such that
-$K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp}.$
-We partition $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ into two subsets, $A_{1}$ and $A_{2}$.
-The set $A_{1}$ consists of all $\xi$ such that, for every $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$,
-$\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}=g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
-for some $g\in G_{K\cap\xi_{0}^{\perp}}$. The set $A_{2}$ consists of all $\xi$ for which there exists a vector $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ such that
-$\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}\neq g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
+First, we introduce two functions $\Lambda:\mathbb{S}^{n-1}\to\mathbb{R}_{\geq 0}$ and $\Gamma:G(n,n-2)\to\mathbb{R}_{\geq 0}$ defined by
+$\Lambda(\theta):=\sup_{\xi,\eta\in\mathbb{S}^{n-1}\cap\theta^{\perp}}\Big\{\Big|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|\Big\}$
+and
+$\Gamma(H):=\sup_{\theta\in\mathbb{S}^{n-1}\cap H\atop\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}}\Big\{\Big|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|\Big\}.$
+It is clear that
+$\Gamma(H)=\sup_{\theta\in\mathbb{S}^{n-1}\cap H}\Lambda(\theta)$
+and Condition (1) is equivalent to $\Lambda(\theta)=0$ for all $\theta\in\mathbb{S}^{n-1}$ or $\Gamma(H)=0$ for all $H\in G(n,n-2)$.
+Moreover, $\Lambda,\Gamma$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$.
+Furthermore, $\Gamma(H_{1})=\Gamma(H_{2})$ whenever there exists a linear transformation between $K\cap H_{1}$ and $K\cap H_{2}$, as implied by Lemma 3.1.
+Now we partition G(n,n2)G(n,n-2) into two subsets, $A_{1}$ and $A_{2}$. The set $A_{1}$ consists of all $H$ such that, for every $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$,
+$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap H)$
+for some $g\in G_{K\cap\xi_{0}^{\perp}}$.
+The set $A_{2}$ consists of all $H\in G(n-2)$ for which there exists a pair of vectors $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$ such that
+$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)\neq\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap H)$
 for every $g\in G_{K\cap\xi_{0}^{\perp}}$. It is clear that $A_{1}\cap A_{2}=\emptyset$.
-We once more partition $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ into two subsets, $B_{1}$ and $B_{2}$. Define
-$B_{1}:=\{\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:\Gamma(\xi)=0\},\quad B_{2}:=\{\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:\Gamma(\xi)>0\}.$
-Our goal is to show that $B_{1}$ and $B_{2}$ are closed.
-Note that, for $\xi\in A_{1}$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
-$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
-where
-$J(K\cap\xi_{0}^{\perp})\cap\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=J(K\cap\xi_{0}^{\perp})\cap g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
+We once more partition G(n,n2)G(n,n-2) into two subsets, $B_{1}$ and $B_{2}$. Define
+$B_{1}:=\{H\in G(n,n-2):\Gamma(H)=0\},\quad B_{2}:=\{H\in G(n,n-2):\Gamma(H)>0\}.$
+Our first goal is to show that if $B_{2}\neq\emptyset$, $B_{2}=G(n,n-2)$.
+Note that, for $H\in A_{1}$ and $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, we have
+$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap H)$
+for some $g\in G_{K\cap\xi_{0}^{\perp}}$. Then
+$\phi_{\eta}(\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp}))=\phi_{\xi}\circ g(\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})),$
+where $g(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, and thus, we have
+$\phi_{\eta}(\phi^{-1}_{\xi}(H)\cap J(K\cap\xi_{0}^{\perp}))=\phi_{\xi}(\phi^{-1}_{\xi}(H)\cap J(K\cap\xi_{0}^{\perp})),$
 which implies
-$J(K\cap\xi_{0}^{\perp})\cap\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(g(J(K\cap\xi_{0}^{\perp}))\cap K\cap\xi^{\perp}\cap\eta^{\perp}),$
-where $g(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, and thus, we have
-$\phi_{\zeta}(J(K\cap\xi_{0}^{\perp})\cap\phi_{\zeta}^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}))=g\circ g^{-1}(J(K\cap\xi_{0}^{\perp})\cap K\cap\xi^{\perp}\cap\eta^{\perp}),$
+$\phi_{\eta}\circ\phi^{-1}_{\xi}(H)\cap\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))=H\cap\phi_{\xi}(J(K\cap\xi_{0}^{\perp})),$
+and thus,
+$\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(H)\cap\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))=H\cap\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$
 that is,
-$\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))\cap\xi_{0}^{\perp}=J(K\cap\xi_{0}^{\perp})\cap\zeta^{\perp}.$
+$\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))\cap H=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap H.$
 Therefore, we obtain
-$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
-for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$. And thus, $\Gamma(\xi)=0$, so $A_{1}\subseteq B_{1}$, which implies $B_{2}\subseteq A_{2}$.
-We now focus on the set $B_{2}$.
-For any $\xi\in B_{2}$, and therefore $\xi\in A_{2}$, and for any $\zeta,\eta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$ and $\zeta\neq\eta$, we have
-$K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp},$
-and, in addition,
-$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\eta^{\perp})=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\zeta}(\eta^{\perp}).$
-Hence, by exploiting the linearity of $\phi^{-1}_{\zeta}$, for any $y\in\phi^{-1}_{\zeta}(\eta^{\perp})$, we obtain $\langle\phi_{\zeta}(y),\eta\rangle=0$.
-Equivalently, this can be written as $y^{T}A^{T}_{\phi_{\zeta}}\eta=0$, which implies that $y\in(A^{T}_{\phi_{\zeta}}\eta)^{\perp}$, where $A_{\phi_{\zeta}}$ denotes the matrix associated with the linear transformation $\phi_{\zeta}$.
-Therefore, there exists a linear transform between $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap(A^{T}_{\phi_{\zeta}}\eta)^{\perp}$.
-Furthermore, the set
-$E_{\zeta,\xi}:=\{A^{T}_{\phi_{\zeta}}\eta:\eta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})\}$
-forms an ellipse, and the set
-$C_{\zeta,\xi}:=\left\{\frac{A^{T}_{\phi_{\zeta}}\eta}{|A^{T}_{\phi_{\zeta}}\eta|}:\eta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})\right\}$
-gives a circle.
-And thus, there exists a linear transform between $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\tilde{\eta}^{\perp}$ for any $\tilde{\eta}\in C_{\zeta,\xi}$.
-Finally, we have
-[…]
-for any $\tilde{\eta}\in C_{\zeta,\xi}$.
-Note that for any other point $\tilde{\xi}\in B_{2}$, there is a circle $C_{\tilde{\zeta},\tilde{\xi}}$ such that
-$\Gamma(\tilde{\xi})=\Gamma(\bar{\eta})$
-for all $\bar{\eta}\in C_{\tilde{\zeta},\tilde{\xi}}$.
-Consequently, the function $\Gamma$ must be constant on $B_{2}$, since any two circles on $\mathbb{S}^{n-1}$ intersect.
-Therefore, $B_{2}$ is closed.
-By the connectedness of $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we obtain either $B_{1}$ or $B_{2}$ as $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$.
-If $B_{1}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we have $\Gamma(\xi)=0$ for any $\xi$, and hence Condition (1) holds.
-If $B_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ holds, then $A_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ implies that any (n2)(n-2)-dimensional sections of $K\cap\xi_{0}^{\perp}$ are all linearly dependent, because for every $\tilde{\eta}\in C_{\zeta,\xi}$ there is a linear transformation mapping $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ onto $K\cap\xi_{0}^{\perp}\cap\tilde{\eta}^{\perp}$.
-Using the fact that all (n1)(n-1)-dimensional sections are linearly dependent, we then conclude that every (n2)(n-2)-dimensional section of $K$ is also linearly dependent.
+$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta),$
+for any $\theta\in\mathbb{S}^{n-1}\cap H$. And thus, $\Gamma(H)=0$, so $A_{1}\subseteq B_{1}$, which implies $B_{2}\subseteq A_{2}$.
+We now focus on the set $B_{2}$. For any $H\in B_{2}$, and therefore $H\in A_{2}$, and for any $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, there exists a linear transformation
+$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})=K\cap\eta^{\perp}\cap\phi_{\eta}\circ\phi^{-1}_{\xi}(\eta^{\perp}).$
+Moreover, since $H\in B_{2}$, there is some $\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ such that
+$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$
+for every $g\in G_{K\cap\xi_{0}^{\perp}}$.
+In particular, this implies $\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq K\cap\xi^{\perp}\cap\eta^{\perp}$, and by the continuity of $\Phi$, there is a path joining $\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$ to $K\cap\xi^{\perp}\cap\eta^{\perp}$ by $K\cap\zeta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ with $\eta$ lying on the arc between $\xi$ and $\eta$.
+Thus, as $\zeta$ moves past $\xi$, there are two possibilities: either $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ returns to $K\cap\xi_{0}^{\perp}\cap\zeta^{\perp}$ exactly $m$ times, giving rise to an $m$-circle, or $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\eta^{\perp})$ arrives at $K\cap\xi_{0}^{\perp}\cap\zeta^{\perp}$, meaning that $\lim_{\zeta\to-\xi}\phi_{\zeta}\circ\phi^{-1}_{\xi}(\theta)=-\theta$ for $\theta\in\xi^{\perp}$.
+Owing to the central symmetry of $K$, we introduce the family $\mathcal{L}(H)$ defined by
+$\mathcal{L}(H):=\{\zeta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp}):\zeta,\xi\in\mathbb{S}^{n-1}\cap H^{\perp}\}.$
+Consequently, $\mathcal{L}(H)\cap\mathbb{S}^{n-1}$ is either a $2m$-circle or a single large circle passing through the antipodal point.
+Moreover, for any $H_{1},H_{2}\in\mathcal{L}(H)\cap G(n,n-2)$, the sections $K\cap H_{1}$ and $K\cap H_{2}$ are linearly dependent.
+We now assume that $B_{2}$ is nonempty, which implies that the maximum value of $\Lambda(\theta)$, denoted by $M$, is nonzero.
+By the definition of $\Gamma$, for every $H\in G(n,n-2)$ containing $\theta$, we have $\Gamma(H)=M$.
+Clearly, such an $H$ lies in $B_{2}$, and thus also in $A_{2}$.
+Hence, by the preceding argument and the fact that $\Gamma$ is invariant under linear transformations, the level set
+$\{H\in G(n,n-2):\Gamma(H)=M\}$
+definitely contains $\mathcal{L}(H)$ for some $H$ satisfying $\Gamma(H)=M$, where $\mathcal{L}(H)$ either contains an (n1)(n-1)-dimensional manifold with antipodal points, or it contains a subset of the form $H\times\infty$, where $\infty$ is the shape of a figure eight.
+Observe that, by the compactness of G(n,n2)G(n,n-2), there exists a sufficiently small $\epsilon>0$ such that for every $c\in(M-\epsilon,M)$, the set
+$\{H\in G(n,n-2):\Gamma(H)=c\}$
+cannot contain $H\times\infty$, and therefore it must contain an (n1)(n-1)-dimensional manifold with antipodal points. Consequently, by the continuity of $\Gamma$, the level set
+$\{H\in G(n,n-2):\Gamma(H)=M\}$
+also contains an (n1)(n-1)-dimensional manifold with antipodal points. This implies that
+$\{H\in G(n,n-2):\Gamma(H)=c\}\mbox{ and }\{H\in G(n,n-2):\Gamma(H)=M\}$
+are the same set, and therefore
+$\{H\in G(n,n-2):\Gamma(H)=M\}=G(n,n-2).$
+We now show that if $B_{2}\neq\emptyset$, then in fact $B_{2}=G(n,n-2)$, which implies $A_{2}=G(n,n-2)$.
+By repeating the preceding argument for $\mathcal{L}(H)$ with any $H\in G(n,n-2)$, we conclude that all (n2)(n-2)-dimensional sections of $K$ are linearly dependent.
 Consequently, by Gromov’s theorem [7], $K$ must be an ellipsoid.
 ∎
-To establish that $E\cap\xi^{\perp}=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$, it suffices to prove that $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ and $\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))$ coincide for any $k$-dimensional subspace ($1\leq k\leq n-2$).
-This leads to the following lemma, which is proved by induction.
-Lemma 3.4.
-Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body.
-Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then one of the following conditions must be satisfied.
-  1. (1)
-    For all $H\in G(n,k)$ with $1\leq k\leq n-2$, we have
-    $\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
-    for every $\theta\in\mathbb{S}^{n-1}\cap H$ and $\xi,\eta\in H^{\perp}$.
-  2. (2)
-    The body $K$ is an ellipsoid.
-Proof.
-If $E$ is not an ellipsoid, then by Lemma 3.3, Condition (1) is satisfied for k=n2k=n-2. Now assume that Condition (1) holds for $k$.
-For any $H\in G(n,k-1)$ and any $\xi\in H^{\perp}$, we have
-$\mathrm{span}(H,\xi)\in G(n,k).$
-Therefore, by the assumption of Condition (1), we obtain
-$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
-for every $\theta\in\mathbb{S}^{n-1}\cap\,\mathrm{span}(H,\xi)$ and every $\zeta,\eta\in\bigl(\mathrm{span}(H,\xi)\bigr)^{\perp}$. In particular, this implies that
-$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
-for all $\theta\in\mathbb{S}^{n-1}\cap H$ and all $\zeta,\eta\in\bigl(\mathrm{span}(H,\xi)\bigr)^{\perp}$, where
-$\dim\bigl((\mathrm{span}(H,\xi))^{\perp}\bigr)=n-1-\dim(H)\geq 2.$
-That Condition (1) is valid for k1k-1 follows from the fact that
-$(\mathrm{span}(H,\xi_{1}))^{\perp}\cap(\mathrm{span}(H,\xi_{2}))^{\perp}\neq\emptyset$
-for any $\xi_{1},\xi_{2}\in H^{\perp}$.
-In particular, in the case $k=1$, for any $\theta\in\mathbb{S}^{n-1}$ and any $\xi,\eta\in\mathbb{S}^{n-1}\cap\theta^{\perp}$, we have
+In particular, Condition (1) in Lemma 3.3 gives
 $\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{E}(\theta).$
@@
 Proof.
-By Lemmas 3.2 and 3.4, either $K$ itself is an ellipsoid with $E=K$, or $E$ is an origin-symmetric star body satisfying $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$.
+By Lemmas 3.2 and 3.3, either $K$ itself is an ellipsoid with $E=K$, or $E$ is an origin-symmetric star body satisfying $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$.
 In the latter case, Theorem 3.1 implies that $E$ is an ellipsoid.

### v5

@@
 Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric star body. If for any $\xi\in\mathbb{S}^{n-1}$, $K\cap\xi^{\perp}$ is a centered ellipsoid, then $K$ is an ellipsoid.
-Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [14, Theorem 3.1]).
+Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [15, Theorem 3.1]).
 The convexity of the body $K$ results from the fact that if we have two points in the body, we can examine the plane they span (with the origin).
@@
 Proof.
-First, we introduce two functions $\Lambda:\mathbb{S}^{n-1}\to\mathbb{R}_{\geq 0}$ and $\Gamma:G(n,n-2)\to\mathbb{R}_{\geq 0}$ defined by
-$\Lambda(\theta):=\sup_{\xi,\eta\in\mathbb{S}^{n-1}\cap\theta^{\perp}}\Big\{\Big|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|\Big\}$
-and
+First, we introduce a function $\Gamma:G(n,n-2)\to\mathbb{R}_{\geq 0}$ defined by
 $\Gamma(H):=\sup_{\theta\in\mathbb{S}^{n-1}\cap H\atop\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}}\Big\{\Big|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|\Big\}.$
-It is clear that
-$\Gamma(H)=\sup_{\theta\in\mathbb{S}^{n-1}\cap H}\Lambda(\theta)$
-and Condition (1) is equivalent to $\Lambda(\theta)=0$ for all $\theta\in\mathbb{S}^{n-1}$ or $\Gamma(H)=0$ for all $H\in G(n,n-2)$.
-Moreover, $\Lambda,\Gamma$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$.
+It is clear that Condition (1) is equivalent to $\Gamma(H)=0$ for all $H\in G(n,n-2)$. And hence, We once more partition G(n,n2)G(n,n-2) into two subsets, $B_{1}$ and $B_{2}$. Define
+$B_{1}:=\{H\in G(n,n-2):\Gamma(H)=0\},\quad B_{2}:=\{H\in G(n,n-2):\Gamma(H)>0\}.$
+Our goal is to show that if $B_{2}\neq\emptyset$, $B_{2}=G(n,n-2)$.
+Moreover, $\Gamma$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$.
 Furthermore, $\Gamma(H_{1})=\Gamma(H_{2})$ whenever there exists a linear transformation between $K\cap H_{1}$ and $K\cap H_{2}$, as implied by Lemma 3.1.
-Now we partition G(n,n2)G(n,n-2) into two subsets, $A_{1}$ and $A_{2}$. The set $A_{1}$ consists of all $H$ such that, for every $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$,
-$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap H)$
+Now we need to deal with the linear transformation inside G(n,n2)G(n,n-2).
+Observe that, for any $H\in G(n,n-2)$ and any $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, there exist two linear maps $\phi^{-1}_{\xi},\phi^{-1}_{\eta}$ that send $H$ back to $\xi_{0}^{\perp}$.
+Depending on the relative positions of $\phi^{-1}_{\xi}(H)$ and $\phi^{-1}_{\eta}(H)$, we divide G(n,n2)G(n,n-2) into two subsets, denoted $A_{1}$ and $A_{2}$.
+The set $A_{1}$ consists of all $H$ such that, for every $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$,
+$\phi^{-1}_{\eta}(K\cap H)=g\circ\phi^{-1}_{\xi}(K\cap H)$
 for some $g\in G_{K\cap\xi_{0}^{\perp}}$.
 The set $A_{2}$ consists of all $H\in G(n-2)$ for which there exists a pair of vectors $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$ such that
-$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)\neq\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap H)$
+$\phi^{-1}_{\eta}(K\cap H)\neq g\circ\phi^{-1}_{\xi}(K\cap H)$
 for every $g\in G_{K\cap\xi_{0}^{\perp}}$. It is clear that $A_{1}\cap A_{2}=\emptyset$.
-We once more partition G(n,n2)G(n,n-2) into two subsets, $B_{1}$ and $B_{2}$. Define
-$B_{1}:=\{H\in G(n,n-2):\Gamma(H)=0\},\quad B_{2}:=\{H\in G(n,n-2):\Gamma(H)>0\}.$
-Our first goal is to show that if $B_{2}\neq\emptyset$, $B_{2}=G(n,n-2)$.
 Note that, for $H\in A_{1}$ and $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, we have
-$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap H)$
+$\phi^{-1}_{\eta}(K\cap H)=g\circ\phi^{-1}_{\xi}(K\cap H)$
 for some $g\in G_{K\cap\xi_{0}^{\perp}}$. Then
-$\phi_{\eta}(\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp}))=\phi_{\xi}\circ g(\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})),$
-where $g(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, and thus, we have
-$\phi_{\eta}(\phi^{-1}_{\xi}(H)\cap J(K\cap\xi_{0}^{\perp}))=\phi_{\xi}(\phi^{-1}_{\xi}(H)\cap J(K\cap\xi_{0}^{\perp})),$
+$\phi^{-1}_{\eta}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})=g\circ\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp}),$
 which implies
-$\phi_{\eta}\circ\phi^{-1}_{\xi}(H)\cap\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))=H\cap\phi_{\xi}(J(K\cap\xi_{0}^{\perp})),$
-and thus,
-$\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(H)\cap\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))=H\cap\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$
+$\phi^{-1}_{\eta}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})=g(\phi^{-1}_{\xi}(K\cap H)\cap g^{-1}(J(K\cap\xi_{0}^{\perp}))),$
+where $g^{-1}(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, and thus, we have
+$\phi^{-1}_{\eta}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})=g(\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})),$
+and hence,
+$\phi^{-1}_{\eta}(K\cap H\cap\phi_{\eta}(J(K\cap\xi_{0}^{\perp})))=g(\phi^{-1}_{\xi}(K\cap H\cap\phi_{\xi}(J(K\cap\xi_{0}^{\perp})))),$
 that is,
-$\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))\cap H=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap H.$
-Therefore, we obtain
+$\phi^{-1}_{\eta}(\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))\cap H)=g\circ\phi^{-1}_{\xi}(\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap H).$
+Therefore, as a result of $\phi^{-1}_{\eta}(K\cap H)=g\circ\phi^{-1}_{\xi}(K\cap H)$, we obtain
+$\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))\cap H=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap H$
+which means
 $\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta),$
@@
 We now focus on the set $B_{2}$. For any $H\in B_{2}$, and therefore $H\in A_{2}$, and for any $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, there exists a linear transformation
-$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})=K\cap\eta^{\perp}\cap\phi_{\eta}\circ\phi^{-1}_{\xi}(\eta^{\perp}).$
-Moreover, since $H\in B_{2}$, there is some $\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ such that
-$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$
+$\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})=K\cap\eta^{\perp}\cap\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(\eta^{\perp}).$
+Moreover, since $H\in A_{2}$, there is some $\eta,\xi\in\mathbb{S}^{n-1}\cap H^{\perp}$ such that
+$\phi^{-1}_{\eta}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$
 for every $g\in G_{K\cap\xi_{0}^{\perp}}$.
-In particular, this implies $\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq K\cap\xi^{\perp}\cap\eta^{\perp}$, and by the continuity of $\Phi$, there is a path joining $\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$ to $K\cap\xi^{\perp}\cap\eta^{\perp}$ by $K\cap\zeta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ with $\eta$ lying on the arc between $\xi$ and $\eta$.
-Thus, as $\zeta$ moves past $\xi$, there are two possibilities: either $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ returns to $K\cap\xi_{0}^{\perp}\cap\zeta^{\perp}$ exactly $m$ times, giving rise to an $m$-circle, or $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\eta^{\perp})$ arrives at $K\cap\xi_{0}^{\perp}\cap\zeta^{\perp}$, meaning that $\lim_{\zeta\to-\xi}\phi_{\zeta}\circ\phi^{-1}_{\xi}(\theta)=-\theta$ for $\theta\in\xi^{\perp}$.
+In particular, we have $\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq K\cap\xi^{\perp}\cap\eta^{\perp}$, and by the continuity of $\Phi$, there is a path joining $\phi_{\eta}\circ g_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$ to $K\cap\xi^{\perp}\cap\eta^{\perp}$ by $K\cap\zeta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ with $\zeta$ lying on the arc between $\xi$ and $\eta$.
+Thus, as $\zeta$ moves past $\xi$, there are two possibilities: either $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ returns to $K\cap\xi^{\perp}\cap\zeta^{\perp}$ exactly $m$ times, giving rise to an $m$-circle, or $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\eta^{\perp})$ arrives at $K\cap\xi^{\perp}\cap\zeta^{\perp}$, meaning that $\lim_{\zeta\to-\xi}\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\theta)=-\theta$ for $\theta\in\xi^{\perp}$.
 Owing to the central symmetry of $K$, we introduce the family $\mathcal{L}(H)$ defined by
-$\mathcal{L}(H):=\{\zeta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp}):\zeta,\xi\in\mathbb{S}^{n-1}\cap H^{\perp}\}.$
-Consequently, $\mathcal{L}(H)\cap\mathbb{S}^{n-1}$ is either a $2m$-circle or a single large circle passing through the antipodal point.
+$\mathcal{L}(H):=\{\zeta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp}):\zeta,\xi\in\mathbb{S}^{n-1}\cap H^{\perp}\}.$
+Consequently, $\mathcal{L}(H)\cap\mathbb{S}^{n-1}$ contains either a $2m$-loop or a single large loop passing through the antipodal point.
 Moreover, for any $H_{1},H_{2}\in\mathcal{L}(H)\cap G(n,n-2)$, the sections $K\cap H_{1}$ and $K\cap H_{2}$ are linearly dependent.
-We now assume that $B_{2}$ is nonempty, which implies that the maximum value of $\Lambda(\theta)$, denoted by $M$, is nonzero.
-By the definition of $\Gamma$, for every $H\in G(n,n-2)$ containing $\theta$, we have $\Gamma(H)=M$.
-Clearly, such an $H$ lies in $B_{2}$, and thus also in $A_{2}$.
+We now assume that $B_{2}$ is nonempty, which implies that the maximum value of $\Gamma(H)$, denoted by $M$, is nonzero.
 Hence, by the preceding argument and the fact that $\Gamma$ is invariant under linear transformations, the level set
@@
 We now show that if $B_{2}\neq\emptyset$, then in fact $B_{2}=G(n,n-2)$, which implies $A_{2}=G(n,n-2)$.
-By repeating the preceding argument for $\mathcal{L}(H)$ with any $H\in G(n,n-2)$, we conclude that all (n2)(n-2)-dimensional sections of $K$ are linearly dependent.
+In other words, for any $H\in G(n,n-2)$, the set $\mathcal{H}$ either contains an (n1)(n-1)-dimensional manifold with antipodal points, or it contains a subset of the form $H\times\infty$, such that for all $H_{1},H_{2}\in\mathcal{L}(H)\cap G(n,n-2)$, the sections $K\cap H_{1}$ and $K\cap H_{2}$ are linearly dependent.
+We now introduce another family $\mathcal{F}(H)$ defined by
+$\mathcal{F}(H):=\{\tilde{H}\in G(n,n-2):K\cap\tilde{H}\ \text{and}\ K\cap H\ \text{are linearly dependent}\}.$
+Observe that $\mathcal{L}(H)\subseteq\mathcal{F}(H)$.
+We then assert that there can be only finitely many pairwise disjoint sets $\{H_{i}\in G(n,n-2)\}_{i=1}^{m}$, each containing a subset of the form $H\times\infty$, such that
+$\mathcal{F}(H_{i})\cap\mathcal{F}(H_{j})=\emptyset$
+for $i\neq j$.
+Suppose, to the contrary, that there are infinitely many pairwise disjoint sets $\{H_{i}\}_{i=1}^{\infty}$.
+Then, for each $i$, the set $H_{i+1}\times\infty$ is contained in $C_{i}$, one of the connected components of the complement determined by $H_{i}\times\infty$.
+By the compactness of G(n,n2)G(n,n-2), both sequences $\{C_{i}\}$ and $\{H_{i}\}$ converge to some limit $\bar{H}$.
+However, for all sufficiently large $i\geq I$, we have
+[…]
+which implies
+$\mathcal{L}(\bar{H})\cap\mathcal{L}(H_{i})\neq\emptyset$
+for all $i\geq I$. This yields a contradiction, since it shows that there must exist a linear transformation relating the $H_{i}$ for all $i\geq I$.
+We now choose $H_{0}$ such that $\mathcal{L}(H_{0})$ contains an (n1)(n-1)-dimensional manifold with antipodal points.
+For any other $H$ for which $\mathcal{L}(H)$ also contains an (n1)(n-1)-dimensional manifold with antipodal points, we obtain
+$\mathcal{L}(H_{0})\cap\mathcal{L}(H)\neq\emptyset,$
+and consequently,
+$\mathcal{L}(H)\subseteq\mathcal{F}(H_{0}).$
+If for some $i\in[1,m]$ we have
+$\mathcal{F}(H_{0})\cap\mathcal{F}(H_{i})\neq\emptyset,$
+then it follows that
+$\mathcal{F}(H_{0})=\mathcal{F}(H_{i}).$
+Otherwise, for every $i\in[1,m]$,
+$\mathcal{F}(H_{0})\cap\mathcal{F}(H_{i})=\emptyset.$
+Consequently, G(n,n2)G(n,n-2) is partitioned into $m$ or $m+1$ pairwise disjoint closed sets, and thus, by the connectedness of G(n,n2)G(n,n-2), we conclude that all (n2)(n-2)-dimensional sections of $K$ are linearly dependent.
 Consequently, by Gromov’s theorem [7], $K$ must be an ellipsoid.
@@
 $\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{E}(\theta).$
-Consequently,
+And thus,
 $E\cap\xi^{\perp}=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi^{\perp}).$

### v8

@@
 Abstract.
-In this paper, we begin by constructing an ellipsoid derived from the John ellipsoid of the sections.
-We then use this framework to present a full proof of Banach’s isometric subspace problem in finite-dimensional spaces via the John ellipsoid.
+In this paper, we begin by constructing two families of origin-symmetric star bodies derived from the John ellipsoid of the sections.
+We then apply the level-set method to obtain a complete proof of Banach’s isometric subspace problem in finite-dimensional spaces via the John ellipsoid.
 Key words and phrases:
@@
 More recently, Ivanov, Mamaev, and Nordskova [8] obtained an affirmative solution for $n=3$.
+This outcome leads from the analysis of sections back to the entire convex body, which provides the original insight for addressing this problem.
 The problem can also be formulated for complex normed spaces; see [4], [7], and [11] for further details.
@@
 Theorem 1.1.
-Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body.
-Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$, if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then $K$ is an ellipsoid.
-The proof of Theorem 1.1 is divided into three steps.
-We first select a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then use the linear invariance property of John ellipsoids to obtain a corresponding star body $E$ inside $K$.
-Secondly, we use the level set method (see [10, 13, 17] for details) to show $E$ an ellipsoid.
-With this large ellipsoid at hand, we then establish the desired conclusion.
+Let $K\subset\mathbb{R}^{n}$ with $n\geq 3$ be an origin-symmetric convex body.
+Fix some direction $\xi_{0}\in\mathbb{S}^{n-1}$.
+Suppose that for every $\xi\in\mathbb{S}^{n-1}$ there exists a linear map $\phi_{\xi}\in GL(n)$ such that
+$K\cap\xi^{\perp}=\phi_{\xi}\bigl(K\cap\xi_{0}^{\perp}\bigr)$
+and moreover $\phi_{\xi}(\xi_{0})=\xi$. Then $K$ must be an ellipsoid.
+Indeed, Banach’s conjecture requires only a linear map on G(n,n1)G(n,n-1). However, to complete the proof, we extend this linear map to the entire space by setting
+$\phi_{\xi}=\varphi_{\xi}\circ\psi_{\xi_{0},\xi},$
+where $\psi_{\xi_{0},\xi}\in SO(n)$ is a rotation sending $\xi_{0}$ to $\xi$, and $\varphi_{\xi}\in GL(n)$ satisfies $\varphi_{\xi}(\xi)=\xi$.
+Then the proof of Theorem 1.1 is divided into three steps.
+We first select a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then use the linear invariance property of John ellipsoids to obtain two families of corresponding star bodies $\{E_{\alpha}\}_{\alpha\in[0,1]}$ and $\{\tilde{E}_{\beta}\}_{\beta\in[0,1]}$ inside $K$.
+Secondly, we apply the level set method (see Süss’s conjecture [10, 13, 17] for further details) to demonstrate that there exists a value $t_{0}$ such that $t_{0}K=E_{t_{0}}=\tilde{E}_{t_{0}}$.
+Having established these identities, we can then derive the desired result.
 2. Notation and Preliminaries
 This section mainly presents some basic content, such as symbols and definitions. For more detailed content, readers can refer to the books by Gardner [6] and Schneider [14].
-Let $\mathbb{R}^{n}$ be the classical $n$-dimensional Euclidean space.
-We denote by $\|x\|$ the Euclidean norm of $x\in\mathbb{R}^{n}$.
-We write $B^{n}=\{x\in\mathbb{R}^{n}:\|x\|\leq 1\}$ for the unit ball and $S^{n-1}=\{x\in\mathbb{R}^{n}:\|x\|=1\}$ for the unit sphere.
-We denote $\{e_{1},e_{2},\dots,e_{n}\}$ to be the standard basis in $\mathbb{R}^{n}$.
-For any $\xi\in S^{n-1}$, we denote $\xi^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle=0\}$ to be a hyperplane that is perpendicular to $\xi$.
-Here, $\langle x,\xi\rangle$ is the usual inner product in $\mathbb{R}^{n}$.
-The Grassmann manifold of $k$-dimensional subspaces in $\mathbb{R}^{n}$ is denoted by $G(n,k)$.
-$GL(n)$ and $SL(n)$ are the general linear group and the special linear group in $\mathbb{R}^{n}$.
-For $\phi\in GL(n)$, we denote $\|\phi\|_{op}=\max_{x\in B^{n}}\|\phi(x)\|$ by the operator norm of $\phi$.
-A set in $\mathbb{R}^{n}$ is called convex if it contains the closed line segment joining any two of its points.
-A convex set is a convex body if it is compact and has a non-empty interior.
-A convex body is strictly convex if its boundary contains no line segments.
-A compact set $L$ is called a star body if the origin $O$ is an interior point of $L$, every line through $O$ meets $L$ in a line segment, and its Minkowski gauge defined by
-$\|x\|_{L}=\inf\{a\geq 0:x\in aL\}$
+Let $\mathbb{R}^{n}$ denote the standard $n$-dimensional Euclidean space.
+For $x\in\mathbb{R}^{n}$, we write $\|x\|$ for its Euclidean norm.
+The unit ball is defined as $B^{n}=\{x\in\mathbb{R}^{n}:\|x\|\leq 1\}$, and the unit sphere as $S^{n-1}=\{x\in\mathbb{R}^{n}:\|x\|=1\}$.
+We use $\{e_{1},e_{2},\dots,e_{n}\}$ to denote the standard basis of $\mathbb{R}^{n}$.
+For any $\xi\in S^{n-1}$, we define
+$\xi^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle=0\}$
+to be the hyperplane orthogonal to $\xi$, where $\langle x,\xi\rangle$ is the usual inner product on $\mathbb{R}^{n}$. Furthermore, we set
+$\xi_{+}^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle\geq 0\}\quad\text{and}\quad\xi_{-}^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle\leq 0\}$
+for the two half-spaces determined by the hyperplane $\xi^{\perp}$.
+The Grassmann manifold of $k$-dimensional subspaces of $\mathbb{R}^{n}$ is written as $G(n,k)$.
+The symbols $GL(n)$ and $SL(n)$ denote, respectively, the general linear group and the special linear group on $\mathbb{R}^{n}$.
+For any $\phi\in GL(n)$, its operator norm is defined by $\|\phi\|_{op}=\max_{x\in B^{n}}\|\phi(x)\|$.
+A subset of $\mathbb{R}^{n}$ is said to be convex if, for any two points in the set, the entire closed line segment joining them lies in the set.
+A convex set is called a convex body if it is compact and has non-empty interior.
+A convex body is strictly convex if no line segment is contained in its boundary.
+A compact set $L$ is called a star body if the origin $O$ lies in the interior of $L$, each line passing through $O$ intersects $L$ in a (possibly degenerate) line segment, and its Minkowski gauge, given by
+$\|x\|_{L}=\inf\{a\geq 0:x\in aL\},$
 is a continuous function on $\mathbb{R}^{n}$.
-[…]
-For $H\in G(n,k)$, we call a body a revolution body with respect to $H$ if, for any special orthonormal transformation $\phi\in SO(n)$ with $\phi(H)=H$, we have $\phi(K)=K$.
+The radial function of $L$ is defined by $\rho_{L}(x)=\|x\|_{L}^{-1}$ for all $x\in\mathbb{R}^{n}\setminus\{O\}$.
+When $x$ lies on the unit sphere $S^{n-1}$, the value $\rho_{L}(x)$ represents the distance from the origin to the boundary of $L$ in the direction of $x$.
 Let $K\subseteq\mathbb{R}^{n}$ be a convex body.
-The John ellipsoid of $K$ is the ellipsoid of maximal volume contained in $K$, denoted by $J(K)$.
-It is clear that the John ellipsoid is invariant under linear transformations; that is, for any $\phi\in GL(n)$, $J(\phi(K))=\phi(J(K))$.
-When the John ellipsoid is the $n$-dimensional unit Euclidean ball $B^{n}$, the body $K$ is said to be in John position.
+The John ellipsoid of $K$, denoted $J(K)$, is defined as the ellipsoid of largest volume that is contained in $K$.
+Clearly, the John ellipsoid behaves equivariantly under linear maps: for every $\phi\in GL(n)$, we have $J(\phi(K))=\phi(J(K))$.
+The level set method (see [10, 13, 17] for more details) consists in identifying a function on the sphere that remains unchanged under linear (congruent) transformations relating sections or projections, so that one of its level sets can coincide with the entire sphere.
+A continuous map $f:X\to Y$ is called null-homotopic if it is homotopic to a constant map.
+In other words, there exists a homotopy $F(x,t):X\times[0,1]\to Y$ such that $F(x,0)=f(x)$ and $F(x,1)$ is a constant map.
+On the sphere $\mathbb{S}^{n-1}$, an equator is an (n2)(n-2)-dimensional subsphere of the form $\mathbb{S}^{n-1}\cap\xi^{\perp}$ for some $\xi\in\mathbb{S}^{n-1}$.
+The geodesic distance between points $\theta_{1}$ and $\theta_{2}$, written $d_{\mathbb{S}^{n-1}}(\theta_{1},\theta_{2})$, is defined as the length of the shorter geodesic segment connecting them, which is equal to the angle between $\theta_{1}$ and $\theta_{2}$.
+The geodesic ball of radius $r$ centered at $o$ is given by
+$\{\theta\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(\theta,o)\leq r\}.$
 3. Proof of the main theorem
@@
 Lemma 3.2.
-Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body.
-Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
-$E:=\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}(J(K\cap\xi_{0}^{\perp})).$
+Let $K\subset\mathbb{R}^{n}$ with $n\geq 3$ be an origin-symmetric convex body.
+Fix some direction $\xi_{0}\in\mathbb{S}^{n-1}$.
+Suppose that for every $\xi\in\mathbb{S}^{n-1}$, there exists a linear map $\phi_{\xi}\in GL(n)$ such that
+$K\cap\xi^{\perp}=\phi_{\xi}\bigl(K\cap\xi_{0}^{\perp}\bigr)$
+and moreover $\phi_{\xi}(\xi_{0})=\xi$.
+Then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ gives rise to two families of origin-symmetric star bodies $\{E_{\alpha}\}_{\alpha\in[0,1]}$ and $\{\tilde{E}_{\beta}\}_{\beta\in[0,1]}$ defined by
+$\begin{cases}E_{\alpha}:=\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cap\alpha(K\cap\xi_{0}^{\perp})\big),\\[4.0pt] \tilde{E}_{\beta}:=\displaystyle\bigcap_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup\beta(K\cap\xi_{0}^{\perp})\big).\end{cases}$
 Proof.
-By the definition of $E$, the set $E$ is an origin-symmetric star set whose radial function is given by
-$\rho_{E}(\theta)=\max\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta):\xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\}.$
-Therefore, to prove that $\rho_{E}(\theta)$ is continuous, it suffices to establish the continuity of the mapping $\xi\mapsto\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$.
+By the definitions of $E_{\alpha}$ and $\tilde{E}_{\beta}$, it follows that both $E_{\alpha}$ and $\tilde{E}_{\beta}$ are origin-symmetric star sets, with radial functions given by
+$\rho_{E_{\alpha}}(\theta)=\sup\big\{\min\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta),\alpha\rho_{K}(\theta)\}:\xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\big\}$
+and
+$\rho_{\tilde{E}_{\beta}}(\theta)=\inf\big\{\max\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta),\beta\rho_{K}(\theta)\}:\xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\big\},$
+respectively, here $\phi_{\xi}(K\cap\xi_{0}^{\perp})=K\cap\xi^{\perp}$.
+Therefore, to prove that $E_{\alpha}$ and $\tilde{E}_{\beta}$ are continuous, it suffices to establish the continuity of the mapping $\xi\mapsto\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$.
 Now we define a set-valued function $\lambda:\mathbb{S}^{n-1}\to GL(n)$ to be
-$\lambda(\xi):=\{\phi_{\xi}\in GL(n):K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})\}.$
+$\lambda(\xi):=\{\phi_{\xi}\in GL(n):K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})\mbox{ and }\phi_{\xi}(\xi_{0})=\xi\}.$
 And we define the linear symmetry group $G_{K\cap\xi_{0}^{\perp}}$ to be
-$G_{K\cap\xi_{0}^{\perp}}:=\{g\in GL(n):g(K\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\}$
+$G_{K\cap\xi_{0}^{\perp}}:=\{g\in GL(n):g(K\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\mbox{ and }g(\xi_{0})=\xi_{0}\}.$
 Note that for any $\phi_{\xi},\psi_{\xi}\in\lambda(\xi)$, we have
@@
 $d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})=0.$
-This implies that $\phi(K\cap\xi_{0})=K\cap\xi^{\perp}$, and hence $\phi\in\Phi(\xi)$. Consequently, we have $\Phi(\xi_{n_{k}})\to\Phi(\xi)$ as $k\to\infty$.
+This implies that $\phi(K\cap\xi_{0})=K\cap\xi^{\perp}$. Moreover,
+$\phi(\xi_{0})=\lim_{k\to\infty}\phi_{\xi_{n_{k}}}(\xi_{0})=\lim_{k\to\infty}\xi_{n_{k}}=\xi.$
+and hence $\phi\in\Phi(\xi)$. Consequently, we have $\Phi(\xi_{n_{k}})\to\Phi(\xi)$ as $k\to\infty$.
 This means that for any subsequence $\{\Phi(\xi_{n_{k}})\}_{k=1}^{\infty}$, there exists a sub-subsequence $\{\Phi(\xi_{n_{k_{l}}})\}_{l=1}^{\infty}$ that converges to $\Phi(\xi)$.
@@
 Remark 3.1.
-The proof of the continuity of $\phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ can also be found in [3, Lemma 1.5].
-Now we consider the origin-symmetric star body $E$. The next step is to confirm that $E\cap\xi^{\perp}$ indeed recovers $\phi_{\xi}(J(E\cap\xi_{0}^{\perp}))$.
+A proof that $\Phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ is continuous is also provided in [3, Lemma 1.5].
+Remark 3.2.
+Observe that the continuous map $\Phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ admits a continuous lift $\phi:\mathbb{S}^{n-1}\to GL(n)$ such that $\phi(\xi_{0})$ is the identity and, for every $\xi$, we can write $\phi(\xi)=\phi_{\xi}\circ g_{\xi}$ with $g_{\xi}\in G_{K\cap\xi_{0}^{\perp}}$.
+Moreover, both $K\cap\xi_{0}^{\perp}$ and $J(K\cap\xi_{0}^{\perp})$ are preserved by $g_{\xi}$.
+Hence, without loss of generality, we will from now on take $\phi_{\xi}$ to be this lift $\phi(\xi)$.
+Our next objective is to define a system $\Gamma$ on $\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}\times[0,1]$, given by
+$\Gamma:=\big\{(\theta,\xi,t)\in\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}\times[0,1]:\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}=t\ \text{and}\ \theta\perp\xi\big\}.$
+Observe that, for any $\xi\in\mathbb{S}^{n-1}$, Lemma 3.1 implies
+$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}$
+$=$ $\big\|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\theta\big\|_{K\cap\xi^{\perp}}$
+$=$ $\big\|\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap\{\lambda\theta:\lambda>0\}\big\|_{K\cap\xi^{\perp}}$
+$=$ $\big\|J(K\cap\xi_{0}^{\perp})\cap\{\lambda\phi^{-1}_{\xi}(\theta):\lambda>0\}\big\|_{\phi^{-1}_{\xi}(K\cap\xi^{\perp})}$
+$=$ $\big\|J(K\cap\xi_{0}^{\perp})\cap\{\lambda\phi^{-1}_{\xi}(\theta):\lambda>0\}\big\|_{K\cap\xi_{0}^{\perp}}$
+$=$ $\rho_{J(K\cap\xi_{0}^{\perp})}(\phi^{-1}_{\xi}(\theta))\big\|\phi^{-1}_{\xi}(\theta)\big\|_{K}.$
+And hence, for any $\eta\in\mathbb{S}^{n-1}$
+$(\theta,\xi,t)\in\Gamma\iff(\phi^{-1}_{\xi}(\theta),\xi_{0},t)\in\Gamma\iff(\phi_{\eta}\circ\phi^{-1}_{\xi}(\theta),\eta,t)\in\Gamma$
+From this point of view, we can construct a continuous mapping $\Lambda_{t}[\theta]:\mathbb{S}^{n-1}\to\mathbb{S}^{n-1}$, defined by
+$\Lambda_{t}[\theta](\eta)=\phi_{\eta}\phi^{-1}_{\xi}(\theta).$ (1)
+Then we can define a level set $\Lambda_{t}$ to be
+$\Lambda_{t}:=\{\theta\in\mathbb{S}^{n-1}:(\theta,\xi,t)\in\Gamma\mbox{ for some }\xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\}.$ (2)
+It is clear that $\Lambda_{t}[\theta](\eta)\subseteq\Lambda_{t}$ for any $\theta\in\Lambda_{t}$. We now present several basic propositions of $\Lambda_{t}[\theta]$ and $\Lambda_{t}$.
+Proposition 3.1.
+Let $K\subset\mathbb{R}^{n}$ with $n\geq 3$ be an origin-symmetric convex body.
+Fix some direction $\xi_{0}\in\mathbb{S}^{n-1}$.
+Suppose that for every $\xi\in\mathbb{S}^{n-1}$, there exists a linear map $\phi_{\xi}\in GL(n)$ such that
+$K\cap\xi^{\perp}=\phi_{\xi}\bigl(K\cap\xi_{0}^{\perp}\bigr)$
+and moreover $\phi_{\xi}(\xi_{0})=\xi$. Then the following propositions hold.
+  1. (1)
+    Either $\Lambda_{t}$ is the entire sphere, or the image of $\Lambda_{t}[\theta]$ is null-homotopic and its complement $\mathbb{S}^{n-1}\setminus\mathrm{Im}(\Lambda_{t}[\theta])$ consists of exactly two connected components;
+  2. (2)
+    Either $\Lambda_{t}$ is the entire sphere, or it is null-homotopic, symmetric with respect to the origin, and contains an equator;
+  3. (3)
+    There exists some $t_{0}\in[0,1]$ for which $\Lambda_{t_{0}}=\mathbb{S}^{n-1}$.
+Before presenting the proof, we note that the $\phi_{\xi}$ appearing here is the continuous lifting introduced in Remark 3.2.
+After this proposition, however, the continuity of $\phi_{\xi}$ is no longer required, since the set $\Lambda_{t}$ does not depend on the continuity of $\phi_{\xi}$.
+Proof.
+First, we aim to prove that if $\Lambda_{t}$ is a proper subset of the sphere, then $\Lambda_{t}[\theta]$ is null-homotopic.
+In fact, if $a\notin\Lambda_{t}$, then $a\notin\Lambda_{t}$ also holds because $K$ is symmetric about the origin.
+Thus, $a,-a\notin\mbox{Im}(\Lambda_{t}[\theta])$, and hence we can define a homotopy
+$F(\eta,s)=\frac{s\,\Lambda_{t}[\theta](\eta)+(1-s)a}{\|s\,\Lambda_{t}[\theta](\eta)+(1-s)a\|}.$
+[…]
+$\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o)\leq r\}.$
+Observe that for any $\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta)\in\Lambda_{t}[\theta]$ and any $\zeta\in\mathbb{S}^{n-1}\cap(\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta))^{\perp}$, the set
+$\{\phi_{\vartheta}\circ\phi^{-1}_{\eta}(\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta)):\vartheta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta),\zeta)\}\subseteq\mathrm{Im}(\Lambda_{t}[\theta])$
+traces out a curve that passes through $\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta)$ and enters both $\mathbb{S}^{n-1}\cap\zeta^{\perp}_{+}$ and $\mathbb{S}^{n-1}\cap\zeta^{\perp}_{-}$ (see Figure 1).
+Figure 1. The level set $\Lambda_{t}[\theta]$ passes through $\zeta^{\perp}$.
+Consequently, every boundary point $x$ of one branch $C$ of $\mathbb{S}^{n-1}\setminus\mathrm{Im}(\Lambda_{t}[\theta])$ must lie on a geodesic ball that is tangent at $x$ and has radius $r_{x}<\frac{\pi}{2}$.
+This subsphere can be written as
+$\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o_{x})\leq r_{x}\}.$
+Otherwise, we may prolong a curve contained in $\mathrm{Im}(\Lambda_{t}[\theta])$ until it intersects $C$.
+By the compactness of $\bar{C}$, this yields a uniform radius $r_{0}<\frac{\pi}{2}$ such that $C$ lies entirely in a geodesic ball of radius $r_{0}$.
+Note that $\mathrm{Im}(\Lambda_{t}[\theta])$ is null-homotopic and $a,-a\notin\mathrm{Im}(\Lambda_{t}[\theta])$.
+Consequently, the complement $\mathbb{S}^{n-1}\setminus\mathrm{Im}(\Lambda_{t}[\theta])$ can only decompose into two connected components, $C_{1}$ and $C_{2}$, which contain $a$ and a-a, respectively.
+Now we prove Proposition (1).
+For (2), assume $\Lambda_{t}$ is not the entire sphere.
+Then for any $\theta\in\Lambda_{t}$, Proposition (1) tells us that $\mathrm{Im}(\Lambda_{t}[\theta])$ is null-homotopic.
+Consequently, for any $\theta_{1},\theta_{2}\in\Lambda_{t}$, we must have
+$\mathrm{Im}(\Lambda_{t}[\theta_{1}])\cap\mathrm{Im}(\Lambda_{t}[\theta_{2}])\neq\emptyset;$
+for if not, then
+$\mathrm{Im}(\Lambda_{t}[\theta_{2}])\subseteq C_{1},$
+[…]
+$\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o_{1})>r_{1}\}$
+for some $r_{1}<\frac{\pi}{2}$, implying that $\tilde{C}_{1}$ contains the hemisphere $\mathbb{S}^{n-1}\cap(o_{1})^{\perp}_{-}$.
+This, however, contradicts the requirement that $\tilde{C}_{1}$ must lie inside a smaller geodesic ball
+$\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o_{1})\leq r_{2}\},$
+for some $r_{2}<\frac{\pi}{2}$. Therefore, $\theta_{1}$ and $\theta_{2}$ are path-connected within $\Lambda_{t}$. Let $\gamma$ be a path in $\Lambda_{t}$ joining them. Then
+$\bigcup_{\theta\in\gamma}\mathrm{Im}(\Lambda_{t}[\theta])$
+itself has to be null-homotopic.
+Moreover, for any $(\theta,\xi,t)\in\Gamma$, the symmetry of $K$ implies that $(-\theta,-\xi,t)\in\Gamma$ as well, which shows that $\Lambda_{t}$ is origin-symmetric.
+Therefore, $\mathbb{S}^{n-1}\setminus\Lambda_{t}$ consists of two antipodal connected components, each of which lies inside one of the geodesic balls
+$\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o)\leq r\}$
+and
+$\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,-o)\leq r\},$
+for some $r<\frac{\pi}{2}$, respectively. Consequently, $\Lambda_{t}$ contains the equator $\mathbb{S}^{n-1}\cap o^{\perp}$. This establishes Proposition (2).
+For (3), observe that there must exist some $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ with $(\theta,\xi_{0},1)\in\Gamma$.
+If this were not the case, then $J(K\cap\xi_{0}^{\perp})$ would lie strictly inside $K\cap\xi_{0}^{\perp}$, contradicting the fact that $J(K\cap\xi_{0}^{\perp})$ is the John ellipsoid.
+If $\Lambda_{1}$ is not equal to the entire sphere, then $\Lambda_{1}$ is null-homotopic and contains some equator.
+For sufficiently small $\epsilon>0$, the union $\bigcup_{t\in(1-\epsilon,1]}\Lambda_{t}$ forms an open set $\mathcal{O}$ that contains $\Lambda_{1}$.
+Assuming that none of the $\{\Lambda_{t}\}_{t\in[1-\epsilon,1]}$ equals the whole sphere, we assert that
+$\Lambda_{1}\subseteq\Lambda_{1-\epsilon}.$
+If this were false, then $\Lambda_{1-\epsilon}$ would provide a null-homotopy that contains $\partial\mathcal{O}$, with $\Lambda_{1}\subseteq\mathcal{O}$ but omitting some points of $\Lambda_{1}$, which is a contradiction.
+Likewise, for any $\epsilon>0$, assuming that none of the sets in the family $\{\Lambda_{t}\}_{t\in[1-\epsilon,1]}$ coincides with the entire sphere, it follows that
+$\Lambda_{q}\subseteq\Lambda_{p}$
+whenever $1-\epsilon\leq p\leq q\leq 1$.
+We define
+$m=\inf_{\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}}\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}.$
+Then any $t$ for which $\Gamma$ holds must lie in the interval $[m,1]$.
+Indeed, if this were not the case, we could find $(\theta,\xi,t)\in\Gamma$ with $t<m$, which would imply $(\phi^{-1}_{\xi}(\theta),\xi_{0},t)\in\Gamma$, contradicting the fact that $m$ is the infimum.
+Thus, we obtain
+$\bigcup_{t\in[m,1]}\Lambda_{t}=\mathbb{S}^{n-1}.$
+Consequently, there must exist some $t_{0}$ such that
+$\Lambda_{t_{0}}=\mathbb{S}^{n-1}.$
+Indeed, if this were not the case, then each $\Lambda_{t}$ for $t\in[m,1]$ would be a proper subset of the sphere. However, by the preceding argument,
+$\Lambda_{m}=\bigcup_{t\in[m,1]}\Lambda_{t}=\mathbb{S}^{n-1},$
+which is a contradiction. ∎
 Lemma 3.3.
-Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body.
-Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then one of the following conditions must be satisfied.
-  1. (1)
-    For all $\xi,\eta\in\mathbb{S}^{n-1}$ we have
-    $\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
-    for every $\theta\in\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\eta^{\perp}$.
-  2. (2)
-    The body $K$ is an ellipsoid.
+Let $K\subset\mathbb{R}^{n}$ with $n\geq 3$ be an origin-symmetric convex body.
+Fix some direction $\xi_{0}\in\mathbb{S}^{n-1}$.
+Suppose that for every $\xi\in\mathbb{S}^{n-1}$, there exists a linear map $\phi_{\xi}\in GL(n)$ such that
+$K\cap\xi^{\perp}=\phi_{\xi}\bigl(K\cap\xi_{0}^{\perp}\bigr)$
+and moreover $\phi_{\xi}(\xi_{0})=\xi$. Then there exists $t_{0}\in[0,1]$ such that
+$t_{0}K=E_{t_{0}}=\tilde{E}_{t_{0}}.$
 Proof.
-First, we introduce a function $\Gamma:G(n,n-2)\to\mathbb{R}_{\geq 0}$ defined by
-$\Gamma(H):=\sup_{\theta\in\mathbb{S}^{n-1}\cap H\atop\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}}\Big\{\Big|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|\Big\}.$
-It is clear that Condition (1) is equivalent to $\Gamma(H)=0$ for all $H\in G(n,n-2)$. And hence, We once more partition G(n,n2)G(n,n-2) into two subsets, $B_{1}$ and $B_{2}$. Define
-$B_{1}:=\{H\in G(n,n-2):\Gamma(H)=0\},\quad B_{2}:=\{H\in G(n,n-2):\Gamma(H)>0\}.$
-Our goal is to show that if $B_{2}\neq\emptyset$, $B_{2}=G(n,n-2)$.
-Moreover, $\Gamma$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$.
-Furthermore, $\Gamma(H_{1})=\Gamma(H_{2})$ whenever there exists a linear transformation between $K\cap H_{1}$ and $K\cap H_{2}$, as implied by Lemma 3.1.
-Now we need to deal with the linear transformation inside G(n,n2)G(n,n-2).
-Observe that, for any $H\in G(n,n-2)$ and any $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, there exist two linear maps $\phi^{-1}_{\xi},\phi^{-1}_{\eta}$ that send $H$ back to $\xi_{0}^{\perp}$.
-Depending on the relative positions of $\phi^{-1}_{\xi}(H)$ and $\phi^{-1}_{\eta}(H)$, we divide G(n,n2)G(n,n-2) into two subsets, denoted $A_{1}$ and $A_{2}$.
-The set $A_{1}$ consists of all $H$ such that, for every $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$,
-$\phi^{-1}_{\eta}(K\cap H)=g\circ\phi^{-1}_{\xi}(K\cap H)$
-for some $g\in G_{K\cap\xi_{0}^{\perp}}$.
-The set $A_{2}$ consists of all $H\in G(n-2)$ for which there exists a pair of vectors $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$ such that
-$\phi^{-1}_{\eta}(K\cap H)\neq g\circ\phi^{-1}_{\xi}(K\cap H)$
-for every $g\in G_{K\cap\xi_{0}^{\perp}}$. It is clear that $A_{1}\cap A_{2}=\emptyset$.
-Note that, for $H\in A_{1}$ and $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, we have
-$\phi^{-1}_{\eta}(K\cap H)=g\circ\phi^{-1}_{\xi}(K\cap H)$
-for some $g\in G_{K\cap\xi_{0}^{\perp}}$. Then
-$\phi^{-1}_{\eta}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})=g\circ\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp}),$
-which implies
-$\phi^{-1}_{\eta}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})=g(\phi^{-1}_{\xi}(K\cap H)\cap g^{-1}(J(K\cap\xi_{0}^{\perp}))),$
-where $g^{-1}(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, and thus, we have
-$\phi^{-1}_{\eta}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})=g(\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})),$
-and hence,
-$\phi^{-1}_{\eta}(K\cap H\cap\phi_{\eta}(J(K\cap\xi_{0}^{\perp})))=g(\phi^{-1}_{\xi}(K\cap H\cap\phi_{\xi}(J(K\cap\xi_{0}^{\perp})))),$
-that is,
-$\phi^{-1}_{\eta}(\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))\cap H)=g\circ\phi^{-1}_{\xi}(\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap H).$
-Therefore, as a result of $\phi^{-1}_{\eta}(K\cap H)=g\circ\phi^{-1}_{\xi}(K\cap H)$, we obtain
-$\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))\cap H=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap H$
-which means
-$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta),$
-for any $\theta\in\mathbb{S}^{n-1}\cap H$. And thus, $\Gamma(H)=0$, so $A_{1}\subseteq B_{1}$, which implies $B_{2}\subseteq A_{2}$.
-We now focus on the set $B_{2}$. For any $H\in B_{2}$, and therefore $H\in A_{2}$, and for any $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, there exists a linear transformation
-$\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})=K\cap\eta^{\perp}\cap\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(\eta^{\perp}).$
-Moreover, since $H\in A_{2}$, there is some $\eta,\xi\in\mathbb{S}^{n-1}\cap H^{\perp}$ such that
-$\phi^{-1}_{\eta}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$
-for every $g\in G_{K\cap\xi_{0}^{\perp}}$.
-In particular, we have $\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq K\cap\xi^{\perp}\cap\eta^{\perp}$, and by the continuity of $\Phi$, there is a path joining $\phi_{\eta}\circ g_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$ to $K\cap\xi^{\perp}\cap\eta^{\perp}$ by $K\cap\zeta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ with $\zeta$ lying on the arc between $\xi$ and $\eta$.
-Thus, as $\zeta$ moves past $\xi$, there are two possibilities: either $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ returns to $K\cap\xi^{\perp}\cap\zeta^{\perp}$ exactly $m$ times, giving rise to an $m$-circle, or $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\eta^{\perp})$ arrives at $K\cap\xi^{\perp}\cap\zeta^{\perp}$, meaning that $\lim_{\zeta\to-\xi}\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\theta)=-\theta$ for $\theta\in\xi^{\perp}$.
-Owing to the central symmetry of $K$, we introduce the family $\mathcal{L}(H)$ defined by
-$\mathcal{L}(H):=\{\zeta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp}):\zeta,\xi\in\mathbb{S}^{n-1}\cap H^{\perp}\}.$
-Consequently, $\mathcal{L}(H)\cap\mathbb{S}^{n-1}$ contains either a $2m$-loop or a single large loop passing through the antipodal point.
-Moreover, for any $H_{1},H_{2}\in\mathcal{L}(H)\cap G(n,n-2)$, the sections $K\cap H_{1}$ and $K\cap H_{2}$ are linearly dependent.
-We now assume that $B_{2}$ is nonempty, which implies that the maximum value of $\Gamma(H)$, denoted by $M$, is nonzero.
-Hence, by the preceding argument and the fact that $\Gamma$ is invariant under linear transformations, the level set
-$\{H\in G(n,n-2):\Gamma(H)=M\}$
-definitely contains $\mathcal{L}(H)$ for some $H$ satisfying $\Gamma(H)=M$, where $\mathcal{L}(H)$ either contains an (n1)(n-1)-dimensional manifold with antipodal points, or it contains a subset of the form $H\times\infty$, where $\infty$ is the shape of a figure eight.
-Observe that, by the compactness of G(n,n2)G(n,n-2), there exists a sufficiently small $\epsilon>0$ such that for every $c\in(M-\epsilon,M)$, the set
-$\{H\in G(n,n-2):\Gamma(H)=c\}$
-cannot contain $H\times\infty$, and therefore it must contain an (n1)(n-1)-dimensional manifold with antipodal points. Consequently, by the continuity of $\Gamma$, the level set
-$\{H\in G(n,n-2):\Gamma(H)=M\}$
-also contains an (n1)(n-1)-dimensional manifold with antipodal points. This implies that
-$\{H\in G(n,n-2):\Gamma(H)=c\}\mbox{ and }\{H\in G(n,n-2):\Gamma(H)=M\}$
-are the same set, and therefore
-$\{H\in G(n,n-2):\Gamma(H)=M\}=G(n,n-2).$
-We now show that if $B_{2}\neq\emptyset$, then in fact $B_{2}=G(n,n-2)$, which implies $A_{2}=G(n,n-2)$.
-In other words, for any $H\in G(n,n-2)$, the set $\mathcal{H}$ either contains an (n1)(n-1)-dimensional manifold with antipodal points, or it contains a subset of the form $H\times\infty$, such that for all $H_{1},H_{2}\in\mathcal{L}(H)\cap G(n,n-2)$, the sections $K\cap H_{1}$ and $K\cap H_{2}$ are linearly dependent.
-We now introduce another family $\mathcal{F}(H)$ defined by
-$\mathcal{F}(H):=\{\tilde{H}\in G(n,n-2):K\cap\tilde{H}\ \text{and}\ K\cap H\ \text{are linearly dependent}\}.$
-Observe that $\mathcal{L}(H)\subseteq\mathcal{F}(H)$.
-We then assert that there can be only finitely many pairwise disjoint sets $\{H_{i}\in G(n,n-2)\}_{i=1}^{m}$, each containing a subset of the form $H\times\infty$, such that
-$\mathcal{F}(H_{i})\cap\mathcal{F}(H_{j})=\emptyset$
-for $i\neq j$.
-Suppose, to the contrary, that there are infinitely many pairwise disjoint sets $\{H_{i}\}_{i=1}^{\infty}$.
-Then, for each $i$, the set $H_{i+1}\times\infty$ is contained in $C_{i}$, one of the connected components of the complement determined by $H_{i}\times\infty$.
-By the compactness of G(n,n2)G(n,n-2), both sequences $\{C_{i}\}$ and $\{H_{i}\}$ converge to some limit $\bar{H}$.
-However, for all sufficiently large $i\geq I$, we have
+By Proposition 3.1 (3), there exists some $t_{0}\in[0,1]$ such that $\Lambda_{t_{0}}=\mathbb{S}^{n-1}$.
+Hence, for every $\theta\in\mathbb{S}^{n-1}$, there is a $\zeta\in\mathbb{S}^{n-1}\cap\theta^{\perp}$ satisfying
+$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}=t_{0}.$
+Consequently,
+$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=t_{0}\rho_{K}(\theta).$
+Using the definitions of $E_{t_{0}}$ and $\tilde{E}_{t_{0}}$, we obtain
+$\rho_{E_{t_{0}}}(\theta)$ $=\sup\big\{\min\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta),t_{0}\rho_{K}(\theta)\}:\ \xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\big\}$
+$=\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=t_{0}\rho_{K}(\theta).$
+Similarly,
+$\rho_{\tilde{E}_{t_{0}}}(\theta)$ $=\inf\big\{\max\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta),t_{0}\rho_{K}(\theta)\}:\ \xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\big\}$
+$=\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=t_{0}\rho_{K}(\theta).$
+Thus, we conclude that
+$t_{0}K=E_{t_{0}}=\tilde{E}_{t_{0}}.$
+Now we can prove the main theorem.
+Proof of the Theorem 1.1.
+By Lemma 3.3, there exists $t_{0}\in[0,1]$ such that
+$t_{0}K=E_{t_{0}}=\tilde{E}_{t_{0}}.$
+On the one hand, we observe that
+$E_{t}$ $=t_{0}K$
+$=\bigcup_{\xi\in\mathbb{S}^{n-1}}t_{0}\phi_{\xi}\big(K\cap\xi_{0}^{\perp}\big)$
+$=\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(t_{0}(K\cap\xi_{0}^{\perp})\big)$
 […]
-which implies
-$\mathcal{L}(\bar{H})\cap\mathcal{L}(H_{i})\neq\emptyset$
-for all $i\geq I$. This yields a contradiction, since it shows that there must exist a linear transformation relating the $H_{i}$ for all $i\geq I$.
-We now choose $H_{0}$ such that $\mathcal{L}(H_{0})$ contains an (n1)(n-1)-dimensional manifold with antipodal points.
-For any other $H$ for which $\mathcal{L}(H)$ also contains an (n1)(n-1)-dimensional manifold with antipodal points, we obtain
-$\mathcal{L}(H_{0})\cap\mathcal{L}(H)\neq\emptyset,$
-and consequently,
-$\mathcal{L}(H)\subseteq\mathcal{F}(H_{0}).$
-If for some $i\in[1,m]$ we have
-$\mathcal{F}(H_{0})\cap\mathcal{F}(H_{i})\neq\emptyset,$
-then it follows that
-$\mathcal{F}(H_{0})=\mathcal{F}(H_{i}).$
-Otherwise, for every $i\in[1,m]$,
-$\mathcal{F}(H_{0})\cap\mathcal{F}(H_{i})=\emptyset.$
-Consequently, G(n,n2)G(n,n-2) is partitioned into $m$ or $m+1$ pairwise disjoint closed sets, and thus, by the connectedness of G(n,n2)G(n,n-2), we conclude that all (n2)(n-2)-dimensional sections of $K$ are linearly dependent.
-Consequently, by Gromov’s theorem [7], $K$ must be an ellipsoid.
-∎
-In particular, Condition (1) in Lemma 3.3 gives
-$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{E}(\theta).$
-And thus,
-$E\cap\xi^{\perp}=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi^{\perp}).$
-Now we can show the following theorem.
-Theorem 3.2.
-Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body.
-Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
-$E:=\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}(J(K\cap\xi_{0}^{\perp})).$
-Moreover, $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$ for any $\xi\in\mathbb{S}^{n-1}$ and $E=J(K)$.
-Proof.
-By Lemmas 3.2 and 3.3, either $K$ itself is an ellipsoid with $E=K$, or $E$ is an origin-symmetric star body satisfying $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$.
-In the latter case, Theorem 3.1 implies that $E$ is an ellipsoid.
-Now take an arbitrary ellipsoid $\tilde{E}$ contained in $K$.
-For every $\xi\in\mathbb{S}^{n-1}$, we have the inequality $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$, because $E\cap\xi^{\perp}$ is the John ellipsoid of the section $K\cap\xi^{\perp}$.
-Since the ellipsoid is an intersection body, the theorem of Koldobsky and Zhang (see [9] and [16]) implies that $|\tilde{E}|\leq|E|$.
-Consequently, $E$ is indeed the John ellipsoid of $K$.
-∎
-Proof of the Theorem 1.1.
-By Theorem 3.2, the body $K$ contains a John ellipsoid $E$ such that, for every $\xi\in\mathbb{S}^{n-1}$, the section $E\cap\xi^{\perp}$ is the John ellipsoid of the section $K\cap\xi^{\perp}$.
-By applying an affine transformation, we can move $K$ into John’s position, denoted by $\tilde{K}$, in which $E$ is transformed into a Euclidean ball $B$.
-Since the John ellipsoid is invariant under linear transformation, the set $\tilde{K}\cap\xi^{\perp}$ has $B\cap\xi^{\perp}$ as its John ellipsoid.
-Consequently, $\phi_{\xi}$ must be a congruence transformation; otherwise, the John ellipsoid of $\phi_{\xi}(\tilde{K}\cap\xi_{0}^{\perp})$ would be $\phi_{\xi}(B\cap\xi^{\perp})$, which differs from $B\cap\xi^{\perp}$.
-Hence, $\tilde{K}\cap\xi^{\perp}$ has the same volume for every $\xi\in\mathbb{S}^{n-1}$, and therefore, by the Minkowski–Funk theorem, $\tilde{K}$ must be a ball, which in turn implies that $K$ is an ellipsoid.
-∎
-4. Acknowledgement
-The crucial idea emerged when the author visited the Tianyuan Mathematical Research Center in Kunming.
+[…]
+[…]
+from which it follows that
+[…]
+On the other hand, we obtain
+$t_{0}K$ $=\tilde{E}_{t}$
+$=\bigcap_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup t_{0}(K\cap\xi_{0}^{\perp})\big)$
+[…]
+[…]
+[…]
+which implies that
+$J(K\cap\xi_{0}^{\perp})\backslash t_{0}(K\cap\xi_{0}^{\perp})=\emptyset.$
+Consequently,
+$J(K\cap\xi_{0}^{\perp})=t_{0}(K\cap\xi_{0}^{\perp}),$
+so $K\cap\xi_{0}^{\perp}$ is an ellipsoid, and by Theorem 3.1 this implies that $K$ itself is an ellipsoid. ∎

### v9

@@
 2. Notation and Preliminaries
-This section mainly presents some basic content, such as symbols and definitions. For more detailed content, readers can refer to the books by Gardner [6] and Schneider [14].
+This section mainly presents some basic content, such as symbols and definitions. For more detailed content, readers can refer to the books by Fuente [6], Gardner [7], and Schneider [14].
 Let $\mathbb{R}^{n}$ denote the standard $n$-dimensional Euclidean space.
@@
 Then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ gives rise to two families of origin-symmetric star bodies $\{E_{\alpha}\}_{\alpha\in[0,1]}$ and $\{\tilde{E}_{\beta}\}_{\beta\in[0,1]}$ defined by
-$\begin{cases}E_{\alpha}:=\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cap\alpha(K\cap\xi_{0}^{\perp})\big),\\[4.0pt] \tilde{E}_{\beta}:=\displaystyle\bigcap_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup\beta(K\cap\xi_{0}^{\perp})\big).\end{cases}$
+$\begin{cases}E_{\alpha}:=\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cap\alpha(K\cap\xi_{0}^{\perp})\big),\\[4.0pt] \tilde{E}_{\beta}:=\left\{\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\big[\xi^{\perp}\backslash\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup\beta(K\cap\xi_{0}^{\perp})\big)\big]\right\}^{c}.\end{cases}$
 Proof.
@@
 respectively, here $\phi_{\xi}(K\cap\xi_{0}^{\perp})=K\cap\xi^{\perp}$.
-Therefore, to prove that $E_{\alpha}$ and $\tilde{E}_{\beta}$ are continuous, it suffices to establish the continuity of the mapping $\xi\mapsto\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$.
+Therefore, to prove that the radial functions of $E_{\alpha}$ and $\tilde{E}_{\beta}$ are continuous, it suffices to establish the continuity of the mapping $\xi\mapsto\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$.
 Now we define a set-valued function $\lambda:\mathbb{S}^{n-1}\to GL(n)$ to be
@@
 A proof that $\Phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ is continuous is also provided in [3, Lemma 1.5].
-Remark 3.2.
-Observe that the continuous map $\Phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ admits a continuous lift $\phi:\mathbb{S}^{n-1}\to GL(n)$ such that $\phi(\xi_{0})$ is the identity and, for every $\xi$, we can write $\phi(\xi)=\phi_{\xi}\circ g_{\xi}$ with $g_{\xi}\in G_{K\cap\xi_{0}^{\perp}}$.
-Moreover, both $K\cap\xi_{0}^{\perp}$ and $J(K\cap\xi_{0}^{\perp})$ are preserved by $g_{\xi}$.
-Hence, without loss of generality, we will from now on take $\phi_{\xi}$ to be this lift $\phi(\xi)$.
+For a fixed $\theta\in\mathbb{S}^{n-1}\cap\xi^{\perp}_{0}$, we introduce the set $\tilde{\Lambda}[\theta]\subseteq\mathbb{S}^{n-1}$ by
+$\tilde{\Lambda}[\theta]:=\{\,\phi_{\xi}(\theta):\phi_{\xi}\in\Phi(\xi)\ \text{and}\ \xi\in\mathbb{S}^{n-1}\,\}.$
+Observe that, for any $\xi\in\mathbb{S}^{n-1}$, Lemma 3.1 implies
+$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\eta)\|\eta\|_{K}$
+$=$ $\big\|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\eta)\eta\big\|_{K\cap\xi^{\perp}}$
+$=$ $\big\|\partial(\phi_{\xi}(J(K\cap\xi_{0}^{\perp})))\cap\{\lambda\eta:\lambda>0\}\big\|_{K\cap\xi^{\perp}}$
+$=$ $\big\|\partial(J(K\cap\xi_{0}^{\perp}))\cap\{\lambda\phi^{-1}_{\xi}(\theta):\lambda>0\}\big\|_{\phi^{-1}_{\xi}(K\cap\xi^{\perp})}$
+$=$ $\big\|\partial(J(K\cap\xi_{0}^{\perp}))\cap\{\lambda\phi^{-1}_{\xi}(\eta):\lambda>0\}\big\|_{K\cap\xi_{0}^{\perp}}$
+$=$ $\rho_{J(K\cap\xi_{0}^{\perp})}(\phi^{-1}_{\xi}(\eta))\big\|\phi^{-1}_{\xi}(\eta)\big\|_{K}.$
+And hence, for any $\eta\in\tilde{\Lambda}[\theta]$, there exists a $\xi\in\mathbb{S}^{n-1}\cap\eta^{\perp}$ such that
+$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\eta)\|\eta\|_{K}=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}$
 Our next objective is to define a system $\Gamma$ on $\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}\times[0,1]$, given by
-$\Gamma:=\big\{(\theta,\xi,t)\in\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}\times[0,1]:\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}=t\ \text{and}\ \theta\perp\xi\big\}.$
-Observe that, for any $\xi\in\mathbb{S}^{n-1}$, Lemma 3.1 implies
-$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}$
-$=$ $\big\|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\theta\big\|_{K\cap\xi^{\perp}}$
-$=$ $\big\|\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap\{\lambda\theta:\lambda>0\}\big\|_{K\cap\xi^{\perp}}$
-$=$ $\big\|J(K\cap\xi_{0}^{\perp})\cap\{\lambda\phi^{-1}_{\xi}(\theta):\lambda>0\}\big\|_{\phi^{-1}_{\xi}(K\cap\xi^{\perp})}$
-$=$ $\big\|J(K\cap\xi_{0}^{\perp})\cap\{\lambda\phi^{-1}_{\xi}(\theta):\lambda>0\}\big\|_{K\cap\xi_{0}^{\perp}}$
-$=$ $\rho_{J(K\cap\xi_{0}^{\perp})}(\phi^{-1}_{\xi}(\theta))\big\|\phi^{-1}_{\xi}(\theta)\big\|_{K}.$
-And hence, for any $\eta\in\mathbb{S}^{n-1}$
-$(\theta,\xi,t)\in\Gamma\iff(\phi^{-1}_{\xi}(\theta),\xi_{0},t)\in\Gamma\iff(\phi_{\eta}\circ\phi^{-1}_{\xi}(\theta),\eta,t)\in\Gamma$
-From this point of view, we can construct a continuous mapping $\Lambda_{t}[\theta]:\mathbb{S}^{n-1}\to\mathbb{S}^{n-1}$, defined by
-$\Lambda_{t}[\theta](\eta)=\phi_{\eta}\phi^{-1}_{\xi}(\theta).$ (1)
+$\Gamma:=\big\{(\eta,\xi,t)\in\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}\times[0,1]:\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\eta)\|\eta\|_{K}=t\ \text{and}\ \eta\perp\xi\big\}.$
 Then we can define a level set $\Lambda_{t}$ to be
-$\Lambda_{t}:=\{\theta\in\mathbb{S}^{n-1}:(\theta,\xi,t)\in\Gamma\mbox{ for some }\xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\}.$ (2)
-It is clear that $\Lambda_{t}[\theta](\eta)\subseteq\Lambda_{t}$ for any $\theta\in\Lambda_{t}$. We now present several basic propositions of $\Lambda_{t}[\theta]$ and $\Lambda_{t}$.
+$\Lambda_{t}:=\{\eta\in\mathbb{S}^{n-1}:(\eta,\xi,t)\in\Gamma\mbox{ for some }\xi\in\mathbb{S}^{n-1}\cap\eta^{\perp}\}.$ (1)
+It is clear that $\tilde{\Lambda}[\theta]\subseteq\Lambda_{t}$ if $\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}=t$.
+Observe that the linear symmetry group is compact ([3, Lemma 2.1]), and therefore has only finitely many connected components.
+We select $H\subseteq G_{K\cap\xi_{0}^{\perp}}$ to be the connected component containing the identity element.
+By the continuity of $\Phi$, for each $\eta\in\Lambda[\theta]$ with $\phi_{\xi}\circ g_{\xi}(\theta)=\eta$, we can choose a local continuous map $\phi_{\varphi}\circ g_{\varphi}:U_{\xi}\to GL(n)$, where $U_{\xi}$ is a neighborhood of $\xi$.
+Fixing $\phi_{\xi_{0}}$ to be the identity, we can construct a continuous path $\gamma$ from $\theta$ to $\eta$ by piecing together these local continuous maps.
+This path yields
+$\phi^{-1}_{\xi}(\eta)\subseteq\{g(\theta):g\in H\}.$
+We redefine
+$\Lambda_{t}[\theta]:=\{$ $\phi_{\xi}(\zeta):\phi_{\xi}\in\Phi(\xi)\mbox{ is local continuous},\ \zeta\in\{g(\theta):g\in H\},$
+$\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}=t,\ \text{and}\ \xi\in\mathbb{S}^{n-1}\}.$
+We now present several basic propositions of $\Lambda_{t}[\theta]$ and $\Lambda_{t}$.
 Proposition 3.1.
@@
   1. (1)
-    Either $\Lambda_{t}$ is the entire sphere, or the image of $\Lambda_{t}[\theta]$ is null-homotopic and its complement $\mathbb{S}^{n-1}\setminus\mathrm{Im}(\Lambda_{t}[\theta])$ consists of exactly two connected components;
+    Either $\Lambda_{t}$ is the entire sphere, or the set of $\Lambda_{t}[\theta]$ is null-homotopic and its complement $\mathbb{S}^{n-1}\setminus\Lambda_{t}[\theta]$ consists of exactly two connected components;
   2. (2)
-    Either $\Lambda_{t}$ is the entire sphere, or it is null-homotopic, symmetric with respect to the origin, and contains an equator;
+    Either $\Lambda_{t}$ is the entire sphere, or $\Lambda_{t}[\theta]\cup\Lambda_{t}[-\theta]$ is null-homotopic, symmetric with respect to the origin, and contains an equator;
   3. (3)
     There exists some $t_{0}\in[0,1]$ for which $\Lambda_{t_{0}}=\mathbb{S}^{n-1}$.
-Before presenting the proof, we note that the $\phi_{\xi}$ appearing here is the continuous lifting introduced in Remark 3.2.
-After this proposition, however, the continuity of $\phi_{\xi}$ is no longer required, since the set $\Lambda_{t}$ does not depend on the continuity of $\phi_{\xi}$.
 Proof.
@@
 In fact, if $a\notin\Lambda_{t}$, then $a\notin\Lambda_{t}$ also holds because $K$ is symmetric about the origin.
-Thus, $a,-a\notin\mbox{Im}(\Lambda_{t}[\theta])$, and hence we can define a homotopy
-$F(\eta,s)=\frac{s\,\Lambda_{t}[\theta](\eta)+(1-s)a}{\|s\,\Lambda_{t}[\theta](\eta)+(1-s)a\|}.$
+Thus, $a,-a\notin\Lambda_{t}[\theta]$, and hence we can define a homotopy
+$F(\xi,g,s)=\frac{s\,\phi_{\xi}\circ g(\theta)+(1-s)a}{\|s\,\phi_{\xi}\circ g(\theta)+(1-s)a\|}.$
 […]
 $\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o)\leq r\}.$
-Observe that for any $\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta)\in\Lambda_{t}[\theta]$ and any $\zeta\in\mathbb{S}^{n-1}\cap(\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta))^{\perp}$, the set
-$\{\phi_{\vartheta}\circ\phi^{-1}_{\eta}(\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta)):\vartheta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta),\zeta)\}\subseteq\mathrm{Im}(\Lambda_{t}[\theta])$
-traces out a curve that passes through $\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta)$ and enters both $\mathbb{S}^{n-1}\cap\zeta^{\perp}_{+}$ and $\mathbb{S}^{n-1}\cap\zeta^{\perp}_{-}$ (see Figure 1).
-Figure 1. The level set $\Lambda_{t}[\theta]$ passes through $\zeta^{\perp}$.
-Consequently, every boundary point $x$ of one branch $C$ of $\mathbb{S}^{n-1}\setminus\mathrm{Im}(\Lambda_{t}[\theta])$ must lie on a geodesic ball that is tangent at $x$ and has radius $r_{x}<\frac{\pi}{2}$.
+Observe that for any $\eta\in\Lambda_{t}[\theta]$ and any $\zeta\in\mathbb{S}^{n-1}\cap\eta^{\perp}$, the set
+$\{\phi_{\vartheta}\circ g_{\vartheta}(\theta)):\vartheta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\eta,\zeta)\}\subseteq\Lambda_{t}[\theta]$
+traces out a curve that passes through $\eta$ and enters both $\mathbb{S}^{n-1}\cap\zeta^{\perp}_{+}$ and $\mathbb{S}^{n-1}\cap\zeta^{\perp}_{-}$ (see Figure 1).
+Figure 1. The set $\Lambda_{t}[\theta]$ passes through $\zeta^{\perp}$.
+Consequently, every boundary point $x$ of one branch $C$ of $\mathbb{S}^{n-1}\setminus\Lambda_{t}[\theta]$ must lie on a geodesic ball that is tangent at $x$ and has radius $r_{x}<\frac{\pi}{2}$.
 This subsphere can be written as
 $\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o_{x})\leq r_{x}\}.$
-Otherwise, we may prolong a curve contained in $\mathrm{Im}(\Lambda_{t}[\theta])$ until it intersects $C$.
+Otherwise, we may prolong a curve contained in $\Lambda_{t}[\theta]$ until it intersects $C$.
 By the compactness of $\bar{C}$, this yields a uniform radius $r_{0}<\frac{\pi}{2}$ such that $C$ lies entirely in a geodesic ball of radius $r_{0}$.
-Note that $\mathrm{Im}(\Lambda_{t}[\theta])$ is null-homotopic and $a,-a\notin\mathrm{Im}(\Lambda_{t}[\theta])$.
-Consequently, the complement $\mathbb{S}^{n-1}\setminus\mathrm{Im}(\Lambda_{t}[\theta])$ can only decompose into two connected components, $C_{1}$ and $C_{2}$, which contain $a$ and a-a, respectively.
+Note that $\Lambda_{t}[\theta]$ is null-homotopic and $a,-a\notin\Lambda_{t}[\theta]$.
+Consequently, the complement $\mathbb{S}^{n-1}\setminus\Lambda_{t}[\theta]$ can only decompose into two connected components, $C_{1}$ and $C_{2}$, which contain $a$ and a-a, respectively.
 Now we prove Proposition (1).
 For (2), assume $\Lambda_{t}$ is not the entire sphere.
-Then for any $\theta\in\Lambda_{t}$, Proposition (1) tells us that $\mathrm{Im}(\Lambda_{t}[\theta])$ is null-homotopic.
-Consequently, for any $\theta_{1},\theta_{2}\in\Lambda_{t}$, we must have
-$\mathrm{Im}(\Lambda_{t}[\theta_{1}])\cap\mathrm{Im}(\Lambda_{t}[\theta_{2}])\neq\emptyset;$
+Then for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we have $\Lambda_{t}[\theta]=-\Lambda_{t}[-\theta]$ by the symmetry of $K$.
+Proposition (1) also tells us that $\Lambda_{t}[\theta]$ and $\Lambda_{t}[-\theta]$ are null-homotopic.
+Consequently, we must have
+$\Lambda_{t}[\theta]\cap\Lambda_{t}[-\theta]\neq\emptyset;$
 for if not, then
-$\mathrm{Im}(\Lambda_{t}[\theta_{2}])\subseteq C_{1},$
-[…]
+$\Lambda_{t}[-\theta]\subseteq C_{1},$
+where $C_{1}$ denotes a connected component of $\mathbb{S}^{n-1}\backslash\Lambda_{t}[\theta]$.
+In this situation, there exists a connected component $\tilde{C}_{1}$ of $\mathbb{S}^{n-1}\backslash\Lambda_{t}[-\theta]$ that contains $\mathbb{S}^{n-1}\backslash C_{1}$; hence $\tilde{C}_{1}$ includes a large geodesic ball of the form
 $\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o_{1})>r_{1}\}$
@@
 $\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o_{1})\leq r_{2}\},$
-for some $r_{2}<\frac{\pi}{2}$. Therefore, $\theta_{1}$ and $\theta_{2}$ are path-connected within $\Lambda_{t}$. Let $\gamma$ be a path in $\Lambda_{t}$ joining them. Then
-$\bigcup_{\theta\in\gamma}\mathrm{Im}(\Lambda_{t}[\theta])$
-itself has to be null-homotopic.
-Moreover, for any $(\theta,\xi,t)\in\Gamma$, the symmetry of $K$ implies that $(-\theta,-\xi,t)\in\Gamma$ as well, which shows that $\Lambda_{t}$ is origin-symmetric.
-Therefore, $\mathbb{S}^{n-1}\setminus\Lambda_{t}$ consists of two antipodal connected components, each of which lies inside one of the geodesic balls
+for some $r_{2}<\frac{\pi}{2}$.
+Therefore, both $\Lambda_{t}[\theta]$ and $\Lambda_{t}[-\theta]$ are null-homotopic with $\Lambda_{t}[\theta]\cap\Lambda_{t}[-\theta]\neq\emptyset$ and $\Lambda_{t}[\theta]=-\Lambda_{t}[-\theta]$.
+And hence, $\Lambda_{t}[\theta]\cup\Lambda_{t}[-\theta]$ is null-homotopy and origin-symmetric.
+Therefore, $\mathbb{S}^{n-1}\setminus(\Lambda_{t}[\theta]\cup\Lambda_{t}[-\theta])$ consists of two antipodal connected components, each of which lies inside one of the geodesic balls
 $\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o)\leq r\}$
@@
 $\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,-o)\leq r\},$
-for some $r<\frac{\pi}{2}$, respectively. Consequently, $\Lambda_{t}$ contains the equator $\mathbb{S}^{n-1}\cap o^{\perp}$. This establishes Proposition (2).
-For (3), observe that there must exist some $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ with $(\theta,\xi_{0},1)\in\Gamma$.
-If this were not the case, then $J(K\cap\xi_{0}^{\perp})$ would lie strictly inside $K\cap\xi_{0}^{\perp}$, contradicting the fact that $J(K\cap\xi_{0}^{\perp})$ is the John ellipsoid.
-If $\Lambda_{1}$ is not equal to the entire sphere, then $\Lambda_{1}$ is null-homotopic and contains some equator.
-For sufficiently small $\epsilon>0$, the union $\bigcup_{t\in(1-\epsilon,1]}\Lambda_{t}$ forms an open set $\mathcal{O}$ that contains $\Lambda_{1}$.
-Assuming that none of the $\{\Lambda_{t}\}_{t\in[1-\epsilon,1]}$ equals the whole sphere, we assert that
-$\Lambda_{1}\subseteq\Lambda_{1-\epsilon}.$
-If this were false, then $\Lambda_{1-\epsilon}$ would provide a null-homotopy that contains $\partial\mathcal{O}$, with $\Lambda_{1}\subseteq\mathcal{O}$ but omitting some points of $\Lambda_{1}$, which is a contradiction.
-Likewise, for any $\epsilon>0$, assuming that none of the sets in the family $\{\Lambda_{t}\}_{t\in[1-\epsilon,1]}$ coincides with the entire sphere, it follows that
-$\Lambda_{q}\subseteq\Lambda_{p}$
-whenever $1-\epsilon\leq p\leq q\leq 1$.
-We define
+for some $r<\frac{\pi}{2}$, respectively. Consequently, $\Lambda_{t}[\theta]\cup\Lambda_{t}[-\theta]$ contains the equator $\mathbb{S}^{n-1}\cap o^{\perp}$. This establishes Proposition (2).
+For (3), we consider the continuous function
+$h(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}.$
+It is evident that $h$ is Lipschitz since both $K\cap\xi_{0}^{\perp}$ and $J(K\cap\xi_{0}^{\perp})$ are convex bodies in $\xi_{0}^{\perp}$.
+We now assume the opposite: that $\Lambda_{t}\neq\mathbb{S}^{n-1}$ for any $t\in[m,1]$, where
 $m=\inf_{\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}}\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}.$
-Then any $t$ for which $\Gamma$ holds must lie in the interval $[m,1]$.
-Indeed, if this were not the case, we could find $(\theta,\xi,t)\in\Gamma$ with $t<m$, which would imply $(\phi^{-1}_{\xi}(\theta),\xi_{0},t)\in\Gamma$, contradicting the fact that $m$ is the infimum.
-Thus, we obtain
-$\bigcup_{t\in[m,1]}\Lambda_{t}=\mathbb{S}^{n-1}.$
-Consequently, there must exist some $t_{0}$ such that
-$\Lambda_{t_{0}}=\mathbb{S}^{n-1}.$
-Indeed, if this were not the case, then each $\Lambda_{t}$ for $t\in[m,1]$ would be a proper subset of the sphere. However, by the preceding argument,
-$\Lambda_{m}=\bigcup_{t\in[m,1]}\Lambda_{t}=\mathbb{S}^{n-1},$
-which is a contradiction. ∎
+Note that for any connected level set of $h$, denoted by $\gamma$, if $\theta\in\gamma$, then
+$\{g(\theta):g\in H\}\subseteq\gamma.$
+Now we choose a connected level set of $h$, denoted by $\gamma$, that divides the sphere into finitely many connected components.
+If, instead, there were infinitely many connected components, then the compactness of $\mathbb{S}^{n-1}$ implies that some of them would have to be arbitrarily small, which would contradict the requirement that each component is an open set.
+Thus, we may assert that there exists at least one connected component $C$ of $\mathbb{S}^{n-1}\setminus\gamma$ such that
+$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}\cup-\overline{C}\ \text{and}\ t\in[m,1]\}=\mathbb{S}^{n-1}.$
+Indeed, if this were not the case, then by Proposition (2), the condition $\cup\{\Lambda_{t}[\theta]:\theta\in\bar{C}\cup-\bar{C}\ \text{and}\ t\in[m,1]\}\neq\mathbb{S}^{n-1}$ would imply that
+$\bigcup\{\Lambda_{t}[\theta]:\theta\in C\cup-C\ \text{and}\ t\in[m,1]\}$
+is an open, null-homotopic subset of $\mathbb{S}^{n-1}$ that contains some equator, whose boundary is contained in
+$\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\}$
+for some $\tilde{t}\in[m,1]$. Consequently, by Proposition (2), we must have
+$\bigcup\{\Lambda_{t}[\theta]:\theta\in C\cup-C\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\};$
+otherwise, the set $\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\}$ would fail to be null-homotopic.
+This shows that, for any connected component $C$ of $\mathbb{S}^{n-1}\setminus\gamma$, we have
+$\bigcup\{\Lambda_{t}[\theta]:\theta\in\bar{C}\cup-\bar{C}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\},$
+and therefore
+$\mathbb{S}^{n-1}=\bigcup\{\Lambda_{t}[\theta]:\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\}\subseteq\Lambda_{\tilde{t}}$,
+which contradicts our assumption.
+We now consider the component $C$ for which
+$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}\cup-\overline{C}\ \text{and}\ t\in[m,1]\}=\mathbb{S}^{n-1}.$
+Then We introduce a partially ordered set $(\mathcal{G},\prec)$ with the following structure: for each $\alpha\in\mathcal{G}$, the curve $\gamma_{\alpha}$ is a connected level set of $\{\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:h(\theta)=t_{\alpha}\}$, and $C_{\alpha}$ denotes one of the connected components of the complement $C\setminus\gamma_{\alpha}$ whose boundary lies on $\gamma_{\alpha}$.
+We then declare $\alpha\prec\beta$ whenever $C_{\alpha}\subsetneq C_{\beta}$.
+observe that if
+$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha}\ \text{and}\ t\in[m,1]\}=\mathbb{S}^{n-1},$
+for any $\beta\succ\alpha$, we have
+$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta}\cup-\overline{C}_{\beta}\ \text{and}\ t\in[m,1]\}=\mathbb{S}^{n-1}.$
+On the other hand, by the previous argument, if
+$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta}\cup-\overline{C}_{\beta}\ \text{and}\ t\in[m,1]\}\neq\mathbb{S}^{n-1},$
+we have
+$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta}\cup-\overline{C}_{\beta}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{t_{\beta}}[\theta]:\theta\in\gamma_{\beta}\cup-\gamma_{\beta}\},$
+hence, for any $\alpha\prec\beta$,
+$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha}\ \text{and}\ t\in[m,1]\}\subseteq\{\Lambda_{t_{\alpha}}[\theta]:\theta\in\gamma_{\alpha}\cup-\gamma_{\alpha}\}$
+$\subseteq$ $\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta}\cup-\overline{C}_{\beta}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{t_{\beta}}[\theta]:\theta\in\gamma_{\beta}\cup-\gamma_{\beta}\}.$
+We now select a totally ordered subset of $\mathcal{G}$, which we denote by $\mathcal{H}$. Define
+$\mathcal{F}:=\Bigl\{\alpha\in\mathcal{H}:\ \bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha},\ t\in[m,1]\}=\mathbb{S}^{n-1}\Bigr\}.$
+We aim to show this set coincides with $\mathcal{H}$, i.e.
+$\mathcal{F}=\mathcal{H}$.
+Otherwise, there exists some $\beta_{0}\in\mathcal{H}$ such that for every $\alpha\prec\beta_{0}$ we have $\alpha\notin\mathcal{F}$, while for every $\beta\succ\beta_{0}$ we have $\beta\in\mathcal{F}$.
+However, for every $\alpha\prec\beta_{0}$ we have
+$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{t_{\alpha}}[\theta]:\theta\in\gamma_{\alpha}\cup-\gamma_{\alpha}\},$
+and therefore
+$\bigcup\{\Lambda_{t}[\theta]:\theta\in C_{\beta_{0}}\cup-C_{\beta_{0}}\ \text{and}\ t\in[m,1]\}$
+$\subseteq$ $\bigcup_{\alpha\prec\beta_{0}}\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha}\ \text{and}\ t\in[m,1]\}$
+$\subseteq$ $\bigcup_{\alpha\prec\beta_{0}}\{\Lambda_{t_{\alpha}}[\theta]:\theta\in\gamma_{\alpha}\cup-\gamma_{\alpha}\}$
+$\subseteq$ $\bigcup\{\Lambda_{t_{\beta_{0}}}[\theta]:\theta\in\gamma_{\beta_{0}}\cup-\gamma_{\beta_{0}}\},$
+which yields
+$\mathbb{S}^{n-1}$ $=\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta_{0}}\cup-\overline{C}_{\beta_{0}}\ \text{and}\ t\in[m,1]\}$
+$\subseteq\bigcup\{\Lambda_{t_{\beta_{0}}}[\theta]:\theta\in\gamma_{\beta_{0}}\cup-\gamma_{\beta_{0}}\}\subseteq\Lambda_{\beta_{0}},$
+contradicting the assumption.
+Now define $\alpha_{0}$ as the infimum of $(\mathcal{H},\prec)$. Clearly, $C_{\alpha_{0}}=\emptyset$, and $\gamma_{\alpha_{0}}$ then satisfies
+$\mathbb{S}^{n-1}$ $=\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha_{0}}\cup-\overline{C}_{\alpha_{0}}\ \text{and}\ t\in[m,1]\}$
+$=\bigcup\{\Lambda_{t_{\alpha_{0}}}[\theta]:\theta\in\gamma_{\alpha_{0}}\cup-\gamma_{\alpha_{0}}\}\subseteq\Lambda_{\alpha_{0}},$
+which contradicts our initial assumption.
+Consequently, our assumption does not hold, and therefore Proposition (3) is valid. ∎
 Lemma 3.3.
@@
 $t_{0}K=$ $\tilde{E}_{t}$
-$=\bigcap_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup t_{0}(K\cap\xi_{0}^{\perp})\big)$
+$=$ $\left\{\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\big[\xi^{\perp}\backslash\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup t_{0}(K\cap\xi_{0}^{\perp})\big)\big]\right\}^{c}$
 […]
+$=$ $\left\{\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\big[\xi^{\perp}\backslash\phi_{\xi}\big(t_{0}(K\cap\xi_{0}^{\perp})\big)\big]\right\}^{c}$
 […]
-[…]
+$=$ $t_{0}K\bigcup\left\{\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\big[\xi^{\perp}\backslash\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\backslash t_{0}(K\cap\xi_{0}^{\perp})\big)\big]\right\}^{c},$
 which implies that